### v1

# A Survey on Point-of-Interest Recommendation: Models, Architectures, and Security

Abstract
The widespread adoption of smartphones and Location-Based Social Networks has led to a massive influx of spatio-temporal data, creating unparalleled opportunities for enhancing Point-of-Interest (POI) recommendation systems. These advanced POI systems are crucial for enriching user experiences, enabling personalized interactions, and optimizing decision-making processes in the digital landscape. However, existing surveys tend to focus on traditional approaches and few of them delve into cutting-edge developments, emerging architectures, as well as security considerations in POI recommendations. To address this gap, our survey stands out by offering a comprehensive, up-to-date review of POI recommendation systems, covering advancements in models, architectures, and security aspects. We systematically examine the transition from traditional models to advanced techniques such as large language models. Additionally, we explore the architectural evolution from centralized to decentralized and federated learning systems, highlighting the improvements in scalability and privacy. Furthermore, we address the increasing importance of security, examining potential vulnerabilities and privacy-preserving approaches. Our taxonomy provides a structured overview of the current state of POI recommendation, while we also identify promising directions for future research in this rapidly advancing field.
Index Terms:
Point-of-Interest Recommendation, Recommender Systems, Large Language Models, Federated Learning
I Introduction
The proliferation of smart devices has fueled the rapid growth of Location-based Social Networks (LBSNs) [1, 2, 3], which allow users to share check-ins, reviews, and personal experiences tied to specific locations. These networks, now with billions of users, generate vast amounts of spatio-temporal data [4, 5, 6, 7, 8, 9], presenting valuable opportunities for personalized Point-of-Interest (POI) recommendations. As a dynamic area in recommendation systems, POI recommendation has gained considerable interest from both users and businesses in recent years. These methods leverage users’ historical check-ins along with multimodal data to suggest personalized destinations [6, 10]. However, the diversity in data size, modality, and user expectations introduces new challenges. These complexities motivate researchers to develop innovative techniques that effectively capture mobility patterns and other relevant features, such as spatial, social, and textual information, to enhance the performance of POI recommendations [11, 12, 13].
POI recommendation research has witnessed significant advancements over the past decade, with researchers continually pushing the boundaries along three dimensions: models, architectures, and security.
  • Model Evolution: From Traditional to Advanced. In the early stages, POI recommendation systems primarily relied on latent factor models like Latent Dirichlet Allocation (LDA) [14] and Matrix Factorization (MF) [15] to model dynamic user behavior [16, 17]. While these methods provided initial solutions, they were limited in capturing the complex patterns of user-POI interactions. The advent of deep learning marked a transformative shift, with models like Long Short-Term Memory (LSTM) networks [18] and the Transformer architecture [19] proving far more capable of capturing global-scale features and dynamic sequences of user behavior. Alongside this deep learning revolution, the exploration of Graph Neural Networks (GNNs) emerged as particularly suitable for learning representations in POI recommendations [20, 21, 22]. GNNs excel at capturing complex dependencies between users and POIs, allowing for more nuanced recommendations. More recently, the field has witnessed rapid advancements with the integration of cutting-edge techniques like Large Language Models (LLMs) [23], Diffusion Models (DMs) [24, 25], and Self-Supervised Learning (SSL) [26]. These innovations have significantly enhanced recommendation accuracy, allowing systems to better model user preferences.
  • Architecture Transformations: From Centralized to Decentralized and Beyond. Initially, POI recommendation systems were predominantly server-side [27, 28], relying on centralized processing to manage data and train models. However, this centralized approach soon faced challenges, particularly with scalability and latency, as the growing demand for real-time recommendations strained system performance. To address these issues, the adoption of edge computing emerged, shifting computation closer to the user’s device. This transition improved responsiveness and real-time capabilities by reducing the dependency on cloud infrastructure. Building on this momentum, recent advancements in federated learning [29] have introduced a decentralized model training approach. By distributing training across multiple devices, federated learning not only enhances system scalability but also offers stronger privacy protections by keeping user data local and reducing the risks of centralized data processing.
  • Security Enhancements: From Vulnerable to Robust and Privacy-Preserving. Parallel to these architectural improvements, POI recommendation systems initially exhibited significant privacy and security vulnerabilities, as early designs were prone to data breaches and exploitation [30, 31, 32]. As these vulnerabilities were exposed, researchers began to focus on developing more secure solutions. Over time, a variety of privacy-preserving techniques were introduced to protect sensitive user data. Approaches such as differential privacy [33] and federated learning [29] have become central to modern POI recommendation systems, ensuring that while user data is safeguarded, the accuracy and relevance of recommendations are maintained. These techniques have shifted the landscape towards more robust systems capable of balancing both security and performance.
While existing surveys [34, 35, 36, 37] have contributed valuable insights into POI recommendation systems, there remains a critical need for a comprehensive review that reflects the rapid developments across POI models, architectures, and security. For instance, while Zhao et al. [37] provide a thorough review of POI recommendation with traditional techniques like matrix factorization, the growing challenges within deep POI recommender systems are not covered. By contrast, Wang et al. [34] provide an overview of various POI recommendation methods in the deep learning era, yet they do not delve deeply into the architectural and security challenges that have emerged with the rise of decentralized systems and practical privacy concerns. Similarly, Islam’s survey [35] emphasizes the impact of deep learning on POI recommendation, but it overlooks key advancements in federated learning and edge computing, which are increasingly shaping the deployment of these systems. Furthermore, Werneck’s survey [36] provides a detailed account of POI recommendation techniques from 2017 to 2019, offering valuable insights into the evolution of methodologies during that time frame, but it lacks coverage of more recent advancements such as the integration of GNNs and LLMs, which bring about not only powerful user modeling capabilities but also intensive computation and scalability issues. To summarize, a noticeable gap in the literature regarding cutting-edge models, architectural evolution, and security considerations is shown in Table I.
TABLE I: In comparing our survey to related surveys in the field, several key distinctions and contributions emerge.
Survey [37] [36] [35] [34] Ours
Taxonomy No Yes Yes Yes Yes
Models Yes Yes Yes Yes Yes
Architecture No No No No Yes
Security No No No No Yes
Year 2016 2020 2022 2023 2024
With the increasing diversity of data sources, the rise of new models and architectures, and the mounting need for privacy-preserving techniques, it is essential to provide an up-to-date, comprehensive survey that covers these critical aspects. By offering an in-depth taxonomy in this article, we aim not only to provide a holistic understanding of POI recommendation but also to pave the way for future research endeavors to build upon this comprehensive groundwork. Our contributions are summarized as follows:
  • We undertake an exhaustive and contemporary assessment of models, architectures, and security facets within POI recommendation systems, providing intricate insights into the varied methodologies and technologies.
  • We not only classify existing research on models, architectures, and security but also introduce a new framework to comprehend and structure these essential elements.
  • Our study highlights several promising areas for future research in POI recommendation, pointing out key topics that are ready for further exploration, encouraging innovation and inspiring researchers to explore new, untapped areas that could shape the future of POI recommendation technology.
The remainder of this paper unfolds as follows: Section II delineates the key concepts within POI recommender systems. Section III presents illustrations of taxonomy. Section IV shows preliminaries of models and studies of them. Section V shows preliminaries of different architectures and existing studies of them. Section VI shows preliminaries of security knowledge and existing studies of them. Section VII outlines potential future directions. Lastly, Section VIII encapsulates the conclusion, highlighting key takeaways.
II Preliminaries
The process of POI recommendation presents a series of challenges centered around identifying suitable upcoming POIs for users based on their historical check-in data and other pertinent information within an LBSN. Consider a group denoted as $U=\left\{u_{1},u_{2},\ldots,u_{N}\right\}$, comprising $N$ users within the LBSN, and a set $P=\left\{p_{1},p_{2},\ldots,p_{M}\right\}$ representing $M$ POIs. Users form connections through a network denoted as $\tilde{U}=\left\{(u_{i},u_{j})|u_{i},u_{j}\in U\right\}$. Each POI $p$ is defined by its geographical coordinates, specifically latitude $x_{p}$ and longitude $y_{p}$, along with a set of attributes $A_{p}$ describing the POI’s semantics. The primary objective is to provide personalized recommendations of relevant POIs based on users’ past interactions and preferences within the LBSN. The key concepts of POI recommendation are articulated as follows:
  • Point-of-Interest: A point of interest, represented as $p_{i}$, signifies a particular location or site recognized for its significance or interest to various individuals, tourists, or researchers. These points on a map hold specific value or importance and can encompass landmarks, attractions, businesses, historical sites, parks, restaurants, hotels, or any locale of potential interest.
  • Check-in Event: A check-in event, denoted as $c^{u}_{o}$, involves a user publicly declaring their presence at a specific venue or location at a specific time step $o$. Typically facilitated through mobile applications or social media platforms, this action entails sharing the user’s current whereabouts with their social circle. During a check-in, the user typically selects or searches for a particular place, such as a restaurant, park, or event space, and indicates their current presence at that spot.
  • Next POI Recommendation: Given a user $u$’s check-in list $C^{u}$, the next POI recommendation task revolves around predicting the subsequent POI at time $o_{i+1}$.
III Taxonomy
This paper distinguishes itself from existing surveys on POI recommendation by adopting a holistic approach that emphasizes three critical and interrelated aspects: models, architectures, and security. This tripartite framework, as illustrated in Fig. 2, serves as a comprehensive taxonomy for systematically analyzing and comparing various POI recommendation systems. To further contextualize the research landscape, we present the distribution of existing studies across these three categories in Fig. 1 (a). The figure underscores a clear trend: methods for improving POI recommendation models dominate the field, comprising over 75% of the total research efforts. This focus reflects the ongoing challenge of enhancing accuracy, personalization, and context-awareness in POI recommendations. On the other hand, studies addressing security issues in POI recommendations remain underrepresented, making up approximately 8% of the existing research. This highlights a significant gap in the literature, given the increasing importance of safeguarding user privacy, defending against adversarial attacks, and ensuring data integrity within location-based services. In contrast, research on architectures, which bridges the gap between models and real-world deployment, occupies the remaining portion, indicating its crucial yet often overlooked role in ensuring the scalability, efficiency, and responsiveness of POI recommendation. In the rest of this section, the introduction of the taxonomy will unfold along these three dimensions.
Refer to caption Refer to caption
(a) Literature distribution (b) Development over time
Figure 1: Overview of POI recommendation studies.
III-A Models
Building on previous POI recommendation surveys [38, 39], which primarily reviewed studies based on traditional non-deep learning and early deep learning methods, there remains a need to address the advances brought by more recent techniques. With the emergence of novel approaches such as diffusion models, self-supervised learning, and large language models, a growing body of research has evolved that is not yet covered in existing surveys. To fill this gap, we provide an extensive review of models spanning from traditional methods to cutting-edge techniques, as illustrated in Fig. 2. To further clarify the progression of these methodologies, we present a visual depiction of their structural evolution in Fig. 1 (b), where the labels are arranged chronologically from bottom to top. This illustration highlights how the research landscape has shifted over time, with an increasing focus on more sophisticated and advanced models. Detailed explanations of each method, along with the corresponding preliminaries, are provided in Section IV.
III-A1 Latent Factor Model-Based
Traditional POI recommendation methods, such as those based on Latent Dirichlet Allocation (LDA) and Matrix Factorization (MF), have played a foundational role in shaping the field, as noted in prior surveys [38, 39]. These approaches generally focus on uncovering latent patterns in user-item interactions to make personalized recommendations. Mathematically, these methods can be formulated as an optimization problem:
$\tilde{V}=\arg\min_{\tilde{V}}\mathcal{L}_{f_{\theta}}(U,\tilde{V})$ (1)
where $f_{\theta}$ represents the function that minimizes the loss function $\mathcal{L}_{f_{\theta}}$. $U$ denotes the list of users for whom POI recommendations are being generated. $\tilde{V}$ represents the list of POIs that minimize the loss function, effectively representing the optimal recommendations.
III-A2 Classic Neural Network-Based
While latent factor models such as LDA and MF have laid the groundwork for POI recommendation, the advent of deep neural networks, particularly LSTM and Transformer models, has introduced a new paradigm. These models significantly improve both recommendation accuracy and the ability to capture contextual and sequential information. By leveraging the strengths of sequential data modeling and complex feature extraction, neural network-based approaches have made substantial contributions to advancing POI recommendation [40, 41]. These models can be formally represented as follows:
$\theta^{\ast}=\arg\min_{\theta}-\sum_{i=1}^{U}\sum_{j=1}^{V}y_{ij}\times\log(p_{ij}(\theta))$ (2)
where $\theta$ denotes the parameters of the neural networks and $U$ is the number of users, $V$ is the number of POIs, $y_{ij}$ represents a binary indicator (1 if user $i$ interacted with POI $j$, 0 otherwise), and $p_{ij}(\theta)$ is the predicted probability of user $i$ interacting with POI $j$, given the model’s parameters $\theta$. This equation represents the cross-entropy loss, which is commonly used in POI recommendation. The goal is to minimize the difference between the predicted probabilities and the actual interaction labels. To differentiate this approach from other deep learning-based methods, we categorize it as classic neural network-based, emphasizing its reliance on traditional neural networks, such as LSTM, GNNs, and Transformers.
III-A3 Self-supervised Learning-Based
Self-supervised learning (SSL) [42] has gained significant attention in the recommendation domain as a powerful approach to address key challenges like data sparsity and lack of labeled data. Traditional POI recommendation systems often rely on large amounts of user-item interaction data, which may be scarce or incomplete in many cases. SSL provides an alternative by leveraging vast amounts of unlabeled data to create pseudo-supervision signals, enabling the model to learn more effective and generalized representations without needing extensive labeled data [43, 44, 45]. This makes SSL particularly well-suited for POI recommendation, where interactions are often sparse, and data about user preferences is limited. In the context of POI recommendation, SSL methods often rely on contrastive learning [43], which creates tasks where the model learns to differentiate between positive and negative examples. This can be formulated as follows:
$\mathcal{L}(\theta)=\sum_{i=1}^{U}\sum_{j=1}^{V}[d(h(u_{i}),h(p_{j}))-d(h(u_{i}),h(p_{j^{\prime}}))]$ (3)
where $\theta$ are the parameters of the self-supervised model. $u_{i}$ is the user embedding for user $i$. $p_{j}$ denotes the POI embedding for POI $j$ (positive example). $p_{j^{\prime}}$ denotes the POI embedding for POI $j^{\prime}$ (negative example). $h(\cdot)$ denotes the embedding function that maps users and POIs to a shared latent space. $d(\cdot)$ denotes the distance function (e.g., Euclidean distance) that measures the similarity between embeddings. In this kind of method, the model minimizes the distance between user embeddings and positive POI embeddings while maximizing the distance between user embeddings and negative POI embeddings. It can also be used to pretrain the recommendation model by encouraging the consistency between augmentations of the same user/item while minimizing similarities between augmentations of different users/items.
III-A4 Generative Model-Based
Generative models use machine learning to discover patterns in data and generate new content. Three commonly used generative models in POI recommendation are large language models (LLMs), diffusion models (DMs), and variational autoencoders (VAEs). LLMs generate coherent text based on input prompts, producing contextually relevant outputs. DMs gradually transform random noise into structured data, creating high-quality images or signals. VAEs encode data into a latent space and decode it to generate new instances.
LLMs, renowned for their ability to comprehend and generate human-like text, have emerged as a transformative force in recommendation systems, including POI recommendations [46, 47]. Traditionally, recommendation models relied on structured data like user-item interactions, but LLMs introduce a paradigm shift by incorporating rich, unstructured textual data, including user reviews, POI descriptions, and contextual insights such as time, location, and trends. This allows LLMs to exceed traditional recommendation approaches constrained by interaction data. By processing this diverse information, LLMs can generate more personalized, context-aware, and nuanced recommendations. Furthermore, LLMs enhance user experience by engaging users in natural dialogue, making interactions more intuitive. The POI recommendation generation process with LLMs can be formalized as follows:
$\mathcal{L}(\theta)=-\sum_{j=1}^{U}\sum_{i=1}^{n}\log p(r_{i}|r_{1},...,r_{i-1},u_{j},c,\theta)$ (4)
where $\theta$ are parameters of the LLMs and adaptation blocks. $u_{j}$ is user profile information. $c$ denotes contextual information (e.g., time and location). $r_{i}$ denotes generated recommendation text. $p(r_{i}|u_{j},c_{i},\theta)$ denotes the probability of generating recommendation $r_{i}$ given user profile $u_{j}$, context $c$, and parameters $\theta$.
DMs, which learn to gradually add noise to data and reverse the process to generate new samples, can be applied to POI recommendation by generating plausible user-POI interaction patterns. They can create new user-POI interaction data, effectively filling in missing entries in the interaction matrix. This addresses data sparsity and improves recommendation accuracy. This research line can be formulated as follows:
$\mathcal{L}(\theta)=E_{x\sim p_{\text{data}}(x)}[D_{KL}(p(x|\theta)||p_{\text{data}}(x))]$ (5)
where $\theta$ denotes parameters of the diffusion model. $x$ is user-POI interaction data (e.g., ratings, visits). $p_{\text{data}(x)}$ is the original data distribution. $p(x|\theta)$ is the distribution of data generated by the diffusion model with parameters $\theta$. $D_{KL}(.)$ is Kullback-Leibler divergence, measuring the difference between two probability distributions.
VAEs [48, 49] are generative models that learn to represent data in a compressed latent space. They consist of an encoder mapping input data to a distribution in this latent space and a decoder reconstructing data from samples drawn from that distribution. This allows VAEs to generate new instances by sampling from the learned latent space. They enhance POI recommendation systems by capturing user preferences and location characteristics in a compressed space. By encoding user interaction data with various POIs, VAEs capture underlying patterns and preferences, generating personalized recommendations and improving user experience in location-based services.
III-B Architectures
The architecture of a POI recommendation system significantly impacts its ability to handle large volumes of data, ensure privacy and security, and deliver real-time, personalized recommendations. To meet the diverse needs of modern applications and address challenges such as data scalability, user privacy, and computational efficiency, POI recommendation architectures have developed into three primary categories: centralized, decentralized-based, and federated learning-based. Detailed explanations of each architecture, along with the corresponding methods, are provided in Section V.
III-B1 Centralized-Based
In traditional centralized-based architectures, user data and POI information are collected, stored, and processed on a central server. These architectures have long been the standard for POI recommender systems [50, 51, 52, 53] due to their ability to aggregate large datasets from multiple users, enabling powerful recommendation algorithms to be applied at scale. However, they face growing concerns related to user privacy, data security, and the cost of managing and scaling centralized infrastructure, particularly as data volumes and user expectations for real-time performance increase. The objective function for optimizing a recommendation model with a centralized architecture can be defined as $\Theta^{*}=\arg\min_{\Theta}\mathcal{L}(\Theta;\mathcal{D})$ where $\Theta$ denotes the model parameters, and $\mathcal{L}(\Theta;\mathcal{D})$ is the loss function calculated over the entire dataset $\mathcal{D}$. This formulation assumes access to all available data in one place, which, while advantageous for performance optimization, amplifies privacy risks and imposes significant computational overhead.
III-B2 Decentralized-Based
Decentralized-based structures shift the processing burden from centralized servers to users’ individual devices (e.g., smartphones) [54]. In this approach, recommendation models are deployed locally on the user’s device, where data processing and model inference occur without the need to continuously communicate with a central server. This architecture offers several advantages for POI recommendation, such as enhanced privacy, since sensitive user data remains on the device, and reduced latency, as recommendations can be generated in real-time without relying on server-side processing. Decentralized approaches are particularly suitable for mobile or offline scenarios, where data privacy is critical, or network connectivity is limited. However, the performance of decentralized models may be constrained by the device’s computational resources, and updating models across devices can be challenging. Formally, the optimization of a decentralized model can be formulated as $\theta^{*}=\arg\min_{\theta}\mathcal{L}(\theta;\mathcal{D}_{u})$, where $\theta$ represents the model parameters specific to the device, and $\mathcal{L}(\theta;\mathcal{D}_{u})$ is the loss function computed over the local dataset $\mathcal{D}_{u}$ belonging to the user. By leveraging local computation, decentralized architectures provide a privacy-preserving solution that enables real-time, personalized recommendations tailored to individual user data.
III-B3 Federated Learning-Based
Federated learning (FL)-based architectures [55] represent a newer paradigm that combines the strengths of both centralized and decentralized approaches. In FL, models are trained collaboratively across multiple user devices, where each device computes updates locally using its own data. These updates are then aggregated by a central server to improve a global model without exposing individual user data. This architecture offers a compelling solution for privacy-preserving POI recommendation, as it keeps user data localized while still benefiting from collective learning. FL also reduces the need for direct data transmission between users and servers, addressing privacy concerns and regulatory requirements (e.g., GDPR). Additionally, it offers scalability and adaptability, allowing models to continuously improve as new data is generated on user devices. However, FL introduces challenges in terms of communication overhead, synchronization, and managing heterogeneous data across devices. In the FL-based architecture, the optimization objective of the global model is defined as $\Theta^{*}=\arg\min_{\Theta}\sum_{k=1}^{K}w_{k}\mathcal{L}_{k}(\Theta)$ where $\Theta$ represents the global model parameters, $K$ is the total number of participating clients, $w_{k}$ is the weight for the $k$-th client, and $\mathcal{L}_{k}(\Theta)$ is the loss function at the $k$-th client.
III-C Security
Security is crucial in designing POI recommendation systems, especially due to the sensitive nature of location data. Recent work has focused on enhancing user privacy and ensuring system security. Techniques like homomorphic encryption and differential privacy are being developed to protect user data while maintaining recommendation accuracy. Additionally, privacy-preserving protocols are implemented to guard against attacks, including tampering with recommendations or misusing user data. These security efforts are closely tied to the models and architectures used in POI systems, ensuring that they provide accurate recommendations while safeguarding user privacy. A secure POI recommendation system must integrate privacy-preserving techniques at both the model and architecture levels, mitigating security threats without compromising performance. Section VI examines the latest research on security in POI recommendations, offering insights into how these considerations are being addressed in practice.
III-C1 Data Integrity Threats
Data integrity threats refer to malicious activities aimed at compromising the accuracy, consistency, and trustworthiness of data within a system. These threats often target the integrity of data at various stages, such as during storage, transmission, or processing. In the context of POI recommendations, data integrity threats can lead to significant issues, such as degraded model performance, biased outputs, and overall system unreliability.
One prominent example of a data integrity threat is poisoning attacks [56, 57], which manipulate the training data in POI recommendation systems. These attacks can significantly affect the reliability and trustworthiness of the system by injecting malicious data that leads to biased outcomes. Formally, the goal of a poisoning attack can be formulated as:
$\tilde{\mathcal{D}}=\arg\max_{\mathcal{D}^{\prime}}\mathcal{L}_{\text{attack}}(\Theta;\mathcal{D}\cup\mathcal{D}^{\prime})$ (6)
where $\tilde{\mathcal{D}}$ represents the manipulated dataset, $\mathcal{L}_{\text{attack}}(\Theta;\mathcal{D}\cup\mathcal{D}^{\prime})$ is the attack loss function, $\Theta$ are the model parameters, and $\mathcal{D}^{\prime}$ is the set of malicious data instances added to the original dataset $\mathcal{D}$.
Additionally, POI recommendation systems may face various other threats, including evasion attacks, where adversaries manipulate input data to deceive the system, and Sybil attacks [58], where multiple fake profiles are created to influence recommendations. The physical trajectory inference attack [59] involves deducing users’ historical trajectories from shared data, potentially compromising their privacy by revealing sensitive location information. To defend against such attacks, robust training methods and anomaly detection techniques are employed to identify and mitigate the impact of poisoned data, ensuring the integrity and performance of the recommendation system.
III-C2 User Privacy Protection
In POI recommendation systems, safeguarding user privacy is paramount, as these systems often handle sensitive personal data. Privacy-enhancing technologies play a crucial role in ensuring that user information remains secure while still enabling meaningful data analysis. Techniques such as decentralized secure computation [60, 61], noise injection [62], and homomorphic encryption [63, 29] are essential in this regard.
For example, homomorphic encryption enables computations on encrypted data, ensuring that raw data remains inaccessible to unauthorized parties. The process can be mathematically expressed as:
$\mathbf{Enc}(x)\oplus\mathbf{Enc}(y)$ $=\mathbf{Enc}(x+y)$ (7)
$\mathbf{Enc}(x)\otimes\mathbf{Enc}(y)$ $=\mathbf{Enc}(x\cdot y)$
where $\mathbf{Enc}(\cdot)$ denotes the encryption algorithm under a specific key, $\oplus$,$\otimes$ , ⊗ represent addition and multiplication over ciphertext, +,+$\cdot$+ , ⋅ represent addition and multiplication over plaintext, and $x,y$ are the input data. This property allows computations to be performed on encrypted data, producing encrypted results that can be decrypted to obtain the final output. These methods facilitate data analysis without compromising the confidentiality of personal information, allowing systems to generate accurate recommendations while maintaining user trust and privacy.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=center, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [POI Recommendation,leaf-head, ver [ Models (Sec. IV), leaf-task,text width=7em [ Latent Factor Model-Based, leaf-task10, text width=11.5em [LDA-Based, leaf-task8, text width=7.5em[LCARS [64, 16], JIM [65], Geo-SAGE [66], TRM
 [65], LA-LDA [67], TRM+ [68], SPORE [69], ST-LDA [70],
UCGT [71], LSARS [72], ST-SAGE [73], TPM [74], modelnode-task8,text width=29em]] [MF-Based, leaf-task9, text width=7.5em[GeoMF [75], GeoMF++ [76], LGLMF [77], SSTPMF [78],
SQPMF [79], modelnode-task9,text width=29em]] ] [ Classic NN-Based, leaf-taska, text width=9.5em [LSTM-Based, leaf-task1, text width=9.5em [ST-LSTM [80], LSTM-S [81], PS-LSTM [82], STMLA [83], LSMA
 [84], AT-LSTM [85], ATSD-GRU [86], DLAN [87],modelnode-task1, text width=29em] ] [Transformer-Based, leaf-task2, text width=9.5em [TLR-M [41], GETNext [88], CARAN [89], JANICP [90],
Li et al. [91], STUIC-SAN [92], AMACF [93], CHA [94],
MVSAN [95], HAT [96], STAR-HiT [97], CAFPR [98],
FPG [99], TGAT [100], MobGT [101], POIBERT [102],
AutoMTN [103], CCDSA [104], Liu  et al. [105], TDGCN [106],
TGAT [100], MGAN [107] , STA-TCN [108] , Xia et al. [109],
STTF-Recommender [110] , BayMAN [111], Kumar et al. [112],
ImNext [113] , BSA-ST-Rec [114] , Wu et al.  [115] , ROTAN [116],modelnode-task2, text width=29em] ] [GNN-Based, leaf-task3, text width=9.5em [STGCN [117], Dynaposgnn [20] , ASGNN [21] , GNN-POI [118] ,
ADQ-GNN [22] , GN-GCN [119] , ATST-GGNN [120] , Zhao et al.
 [121] , HS-GAT [122] , PROG-HGNN [28] , CGNN-PRRG [123] ,
HKGNN [124],modelnode-task3, text width=29em] ] ] [ SSL-Based Models, leaf-task6, text width=9.5em [SML [125], S2GRec [126], SSTGL [127], SelfTrip [128], GSBPL [44], LSPSL [129],
Gao et al. [130], Wang et al. [131], SLS-REC [132], CLLP [45], , modelnode-task6, text width=40em] ] [ Generative-Based, leaf-taska, text width=9.5em [ LLM-Based, leaf-task5, text width=9.5em [LLMmove [46], Li et al. [47], modelnode-task5,text width=29em] ] [ DM-Based, leaf-task7, text width=9.5em [Diff-POI [133], DSDRec [134] , Diff-DGMM [135] , DCPR [136], modelnode-task7,text width=29em] ] [VAE-Based, leaf-task4, text width=9.5em [WaPOIR [137], Li et al. [138], modelnode-task4, text width=29em] ] ] ] [ Architectures
(Sec. V), leaf-paradigms,text width=7em [ Centralized, leaf-paradigms, text width=9.5em [ GeoIE [51], SH-CDL [50], GETnext [88], STGN [139], AGRAN [53] , modelnode-paradigms, text width=40em] ] [ Decentralized-Based, leaf-paradigms, text width=9.5em [ MAC [54], DARD [140], LLRec [141], DCLR [27], modelnode-paradigms, text width=40em] ] [ FL-Based, leaf-paradigms, text width=9.5em [ FL&PP-POI [142], Huang et al. [143],CPF-POI [144], SCFL [145], SFL [146], FedPoiRec [29],
PrefFedPOI [147], PREFER [148], Dong et al. [149], DMF [60], NRDL [150] , modelnode-paradigms, text width=40em] ] ] [ Security (Sec. VI), leaf-others,text width=7em [ Data Integrity Threats, leaf-others, text width=9.5em [FedRecAttack [56], Ghost riders [58] , LOKI [57], Yuan et al. [151], PTIA [59] , modelnode-others, text width=40em] ] [ User Privacy Protection, leaf-others, text width=9.5em [DMF [152], Li et al. [62], PriRec [61], Liu et al. [63], Huo et al. [153], FedPoiRec [29], modelnode-others, text width=40em] ] ] ]
Figure 2: Taxonomy of existing studies for POI recommendation in terms of models, architecture and security
IV Models
This section provides a preliminary overview of the models relevant to our survey. We also review specific studies in terms of published venues, techniques, sub-tasks, and datasets in Table LABEL:tab:gen_method.
IV-A Latent Factor Model-Based
IV-A1 Latent Dirichlet Allocation
Early studies in POI recommendation leveraged the power of topic modeling, specifically latent Dirichlet allocation (LDA) [14], to uncover hidden thematic structures within user activity data. LDA, a generative probabilistic model, assumes that POIs are composed of a mixture of latent topics, each characterized by a distribution over words. This approach, while effective in identifying general themes, often lacked the granularity to capture individual preferences and the dynamic nature of user behaviors.
To address these limitations, researchers began exploring more personalized and context-aware approaches [72, 74, 71, 73, 68, 69, 154, 70, 67, 16, 65, 66]. A pioneering work by Yin et al. [16] introduced LCARS, a location-content-aware POI recommender system that combined individual user preferences with the distinct characteristics of a location. This marked a significant shift from purely topic-based models by explicitly incorporating the influence of local trends and preferences on user choices. Building upon this foundation, Wang et al. [74] proposed the Temporal Personalized Model (TPM), which further enhanced POI recommendation by incorporating temporal dynamics and user-specific behavior patterns. TPM introduces the novel concept of topic-regions, clustering locations into semantically meaningful groups based on user activity.
IV-A2 Matrix Factorization
Matrix factorization (MF) [15] is a powerful technique for POI recommendation, addressing sparse user-item interaction data by decomposing it into low-dimensional matrices representing latent user and POI features. This reveals hidden patterns and preferences, enabling personalized recommendations by predicting user affinity for unvisited POIs. Popular methods [154, 12, 75, 76] include SVD [155], PMF [156], and NMF [157]. Specifically, users tend to exhibit spatially clustered mobility patterns, with visits concentrated within specific geographical regions. This “clustering phenomenon” has proven valuable in enhancing POI recommendations, prompting its integration into factorization models. To leverage this insight, Lian et al. [75] propose that user representations are enriched with “activity area vectors” capturing their typical movement zones, while POI representations incorporate “influence area vectors” reflecting the regions from which they draw visitors. This fusion of spatial clustering with factorization models allows for a more context-rich and accurate representation of user preferences and POI characteristics, ultimately leading to more effective personalized POI recommendations.
IV-B Classic Neural Network-Based
IV-B1 Long Short-Term Memory
The long short-term memory (LSTM) network [18] is a specialized type of sequential neural network that allows information to persist over time. It is an enhancement of the recurrent neural network (RNN) [158], specifically designed to address the vanishing gradient problem, a common issue in RNNs. In the original LSTM architecture [18], at each time step $t$, the input $x_{t}$ is combined with the previous hidden state $h_{t-1}$. This combined vector is then processed through three key components: the input gate $\mathsf{input}_{t}$, the output gate $\mathsf{output}_{t}$, and the input node $\mathsf{gate}_{t}$, defined by the following equations:
$\mathsf{input}_{t}$ $=\phi(\mathbf{W}^{(i)}x_{t}+\mathbf{U}^{(i)}h_{t-1}+b^{(i)})$ (8)
$\mathsf{output}_{t}$ $=\phi(\mathbf{W}^{(o)}x_{t}+\mathbf{U}^{(o)}h_{t-1}+b^{(o)})$
$\mathsf{gate}_{t}$ $=\mathsf{tanh}(\mathbf{W}^{(g)}x_{t}+\mathbf{U}^{(g)}h_{t-1}+b^{(g)})=\widetilde{C}_{t}$
where $\mathbf{W}^{(*)}$ and $\mathbf{U}^{(*)}$ are weight matrices, $b^{(*)}$ are bias terms, $\widetilde{C}_{t}$ is the candidate cell state, and $\phi$ is the activation function.
LSTM networks have become indispensable for POI recommendation [80, 52, 81, 83, 84], effectively modeling the sequential dependencies inherent in user check-in data to predict future preferences. Earlier studies [80, 52, 81] adopt LSTM to model POI sequences. Zhang et al. [80] propose a novel approach to learning user preferences by introducing the Discrete Ranking-based Matrix Factorization (DRMF) algorithm. DRMF leverages pairwise user preferences, framing the learning process as a binary quadratic programming problem to derive binary codes. To address the challenges of non-convexity and binary constraints inherent in this formulation, they introduce a self-paced learning strategy. This strategy gradually incorporates pairwise preferences, starting with simpler relationships and progressing to more complex ones, thereby enhancing the optimization process. Nevertheless, Zhao et al. [52] address the limitations of traditional RNNs in capturing the crucial influence of time and distance between consecutive check-ins. Their proposed STLSTM model integrates spatio-temporal context, short-term spatio-temporal gating, and long-term spatio-temporal gating. Another earlier study [81] introduces an approach that incorporates temporal information to understand the semantic connections between user actions. This approach leverages a semantic correlational graph to calculate semantic sequential correlation, factoring in time intervals. It then integrates this semantic information into a modified LSTM framework, incorporating two additional semantic gates. This enhanced model captures users’ sequential behaviors and long- and short-term interests within a semantic context. Finally, this approach proposes user clustering to improve recommendation accuracy by grouping users based on their semantic preferences.
Recent studies [84, 87] based on LSTM mainly aim to capture dynamic dependencies in long-term POI sequences. Yang et al. [84] tackle the challenge of effectively capturing the dynamic interplay between a user’s long-term preferences and their more immediate, context-dependent interests, namely AMACF. It achieves this by dynamically weighting input sequences and seamlessly fusing long- and short-term signals. DLAN [87] incorporates a multi-head attention module that effectively combines first-order and higher-order neighborhood information within user check-in trajectories. This parallel approach addresses the limitations of RNN-based methods, which struggle to establish long-term dependencies between sequences. Furthermore, DLAN integrates a user similarity weighting layer to quantify the influence of social relationships between users on the target user’s preferences. This layer leverages social connections to improve the accuracy of recommendations by incorporating insights from similar users.
Beyond these examples, other studies such as Zhong et al. [82] and Zhang et al. [83] also explore innovative ways to leverage LSTMs for POI recommendation, further demonstrating the versatility and effectiveness of this approach. These studies highlight the ongoing evolution of LSTM-based methods for POI recommendation. Researchers continue to develop sophisticated techniques to capture the complexities of user behavior, ultimately leading to more accurate, personalized, and context-aware recommendations.
IV-B2 Transformer
The Transformer architecture [19], originally designed for natural language processing, is a deep neural network model that heavily relies on the self-attention mechanism. The architecture comprises two main components: an encoder and a decoder, each consisting of multiple identical layers known as transformer blocks. The encoder’s primary function is to process input data and generate corresponding encodings, while the decoder utilizes these encodings, along with contextual information, to generate the final output sequence. As shown in Fig. 3, each transformer block includes several key components: a multi-head attention mechanism, a feed-forward neural network and residual connections [39].
Figure 3: The Transformer model architecture
Transformers have revolutionized POI recommendation systems [41, 88, 89, 90, 91, 100], surpassing traditional methods by effectively capturing complex sequential patterns and contextual information crucial for understanding user preferences. A former study [41] addresses the challenge of providing queue-time-aware POI recommendations. This task is non-trivial, as it requires both recommending the next POI and accurately predicting the queue time at that location. To tackle this challenge, Halder et al. [41] propose TLR-M, a multi-task, multi-head attention transformer model. TLR-M simultaneously recommends POIs to target users and predicts the queue time for accessing those POIs. The model’s multi-head attention mechanism efficiently integrates long-range dependencies between any two POI visits, enabling it to assess their influence on POI selection.
A recent study [88] addresses the limitations of treating POI recommendation solely as a sequence prediction task by incorporating collaborative signals through the proposed method GETNext. GETNext introduces a user-agnostic map to capture global movement patterns and collaborative information. It also integrates the flow map into the transformer architecture, enabling the model to learn from both individual user sequences and collective user behavior. Meanwhile, leveraging collaborative signals proves particularly beneficial for recommending POIs to new users with limited historical data. Explicit user preference integration [91] focuses on directly incorporating explicit user preference data into the transformer model, contrasting with other methods that rely solely on implicit preference learning from POI sequences. It explicitly incorporates data reflecting user preferences, such as ratings or reviews, into the model’s input. Directly modeling user preferences aims to provide a more nuanced and accurate representation compared to solely inferring preferences from interaction patterns. The Temporal-Geographical Attention-based Transformer [100] tackles the challenge of effectively integrating multiple contextual factors throughout the POI recommendation process. Unlike previous studies, TGAT dynamically selects POI sequences from multiple contextual factor POI graphs derived from user check-in histories to obtain better user representations. Then, TGAT incorporates diverse contextual factors, such as time of day, user interests, and geographical influences, throughout the representation learning process. TGAT also integrates contextual factors, leading to more holistic and accurate representations of POIs and users, thereby enhancing recommendation accuracy.
IV-B3 Graph Neural Networks
Graph neural networks (GNNs) [159] provide a powerful framework for learning representations from graph-structured data. GNNs operate through a message-passing mechanism, where each node updates its representation by aggregating information from its neighboring nodes and then combining these aggregated messages with its current representation. An illustration of the GNN architecture is presented in Fig. 4 (a).
Consider a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. Each node $v\in V$ is associated with a feature vector $h^{\prime}_{v}$. GNNs utilize the structure of the graph, along with these node features $h^{\prime}_{v}$, to learn a representation vector for each node, $h_{v}$, or for the entire graph, $h_{G}$. Modern GNN architectures typically employ a neighborhood aggregation strategy, where a node’s representation is iteratively refined by aggregating the representations of its neighbors [160]. After $k$ iterations of aggregation, the node’s representation captures the structural information within its $k$-hop neighborhood. The $k$-th layer of a GNN is shown as:
$a_{v}^{(k)}$ $=\mathsf{Aggregate}^{(k)}(\{h_{u}^{(k-1)}:u\in\mathcal{N}(v)\})$ (9)
$h_{v}^{(k)}$ $=\mathsf{Combine}^{(k)}(h_{v}^{(k-1)},a_{v}^{(k)})$
where $h_{v}^{(k)}$ is the feature vector of node $v$ at the $k$-th iteration, with $h_{v}^{(0)}$ initialized to $X_{v}$. The set $\mathcal{N}(v)$ denotes the nodes adjacent to $v$. The functions $\mathsf{Aggregate}^{(k)}(\cdot)$ and $\mathsf{Combine}^{(k)}(\cdot)$ are crucial to the design of GNNs.
GNNs have emerged as a powerful tool for POI recommendation [117, 20, 21, 118, 120, 123, 124], effectively capturing complex relationships within location-based data. Earlier studies [117, 20, 21, 118, 120] mainly emphasize capturing spatial and temporal dynamics. Specifically, Han et al. [117] focuses on capturing both spatial and temporal dynamics within user check-in data. First, it constructs a multigraph to represent various contextual factors and their relationships. Then, it employs a time-aware sampling strategy to capture the evolving influence of neighboring nodes. Finally, it learns node representations that change over time, reflecting the dynamic nature of user preferences and POI popularity. Another earlier study [20] proposes a novel next POI recommendation model called DynaPosGNN, which incorporates specific arrival times into offline user activities. Unlike traditional methods that solely rely on user check-in history, DynaPosGNN analyzes the correlation between arrival time and two spatial dynamic graphs: the “User-POI graph” and the “POI-POI graph”. By integrating temporal information, DynaPosGNN provides a more nuanced understanding of user behavior, considering both the locations visited and the timing of those visits, leading to potentially more accurate recommendations. In comparison, ASGNN [21] incorporates a personalized hierarchical attention network to exploit complex correlations between users and POIs in check-in sequences. This allows the model to capture both long-term and short-term user preferences, providing a more discriminative understanding of user behavior. Finally, ASGNN leverages these learned user preferences to predict the next POI, taking into account both long-term and short-term behavioral patterns. Zhang et al. [118] introduce GNN-POI, a POI recommendation framework that leverages the power of GNNs to learn node representations from both node information and topological structure, thus enhancing POI recommendation accuracy. GNNs have demonstrated exceptional capacity for modeling complex relationships within graph structures, making them well-suited for capturing the intricate connections between users and POIs in LBSNs.
Refer to caption Refer to caption
(a) GNN model structure (b) DM model sturcture
Figure 4: GNN and DM model structures.
Recent studies [124, 123] based on GNN focus on adopting GNN to capture side information. Liu et al. [123] propose an approach to learning similar users’ POI transfer preferences using a session-aware GNN. This method aims to understand how users’ preferences for POIs evolve over time and how those preferences can be transferred to similar users. By contrast, HKGNN [124] leverages the power of hypergraphs and side information to enhance POI recommendations. The key details include utilizing a hypergraph to represent the complex relationships between users, POIs, and other relevant entities, employing a hypergraph neural network to learn from the rich structural information encoded in the HKG, and incorporating side information, such as POI categories and user demographics, to mitigate data sparsity and enhance recommendation accuracy.
IV-C Self-Supervised Learning-Based
Self-supervised learning (SSL) [42] is a versatile framework that leverages unsupervised data by creating surrogate tasks, known as pretext tasks, to help models learn meaningful representations. These tasks are specifically designed to guide the learning process towards capturing useful features from the data. SSL methods are generally categorized into two approaches: generative and contrastive [161]. Generative SSL aims to create or reconstruct data, focusing on learning detailed representations, while contrastive SSL emphasizes comparing data samples by bringing similar pairs closer in the latent space and pushing dissimilar pairs apart.
The SSL process typically begins with defining a pretext task, which does not require labeled data but helps in learning valuable representations. In contrastive learning, the pretext task might involve differentiating between augmented versions of the same image (positive pairs) and different images (negative pairs). The model is trained by minimizing the following loss function [162]:
$\mathcal{L}_{\text{SSL}}=-\log\frac{\mathsf{Exp}(\mathsf{Sim}(\mathsf{pair}_{i},\mathsf{pair}_{j})/\tau)}{\sum_{k=1}^{2N}\mathbf{1}_{[k\neq i]}\mathsf{Exp}(\mathsf{Sim}(\mathsf{pair}_{i},\mathsf{pair}_{k})/\tau)}$ (10)
where $\mathbf{1}_{[k\neq i]}\in\{0,1\}$ is an indicator function that equals 1 only if $k\neq i$, $\mathsf{pair}_{i}$ and $\mathsf{pair}_{j}$ are the feature representations of a positive pair, $\mathsf{Sim}(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is a temperature scaling parameter, and $N$ represents the number of data points.
Self-supervised learning has emerged as a promising paradigm in POI recommendation [163]. By leveraging patterns and structures within data, self-supervised techniques enable models to learn meaningful representations without explicit labels. In POI recommendation, these methods capture latent user preferences and spatial contexts from unlabeled data, enhancing personalized recommendations. Self-supervised POI recommendation systems [125, 126, 128, 129, 44, 131, 132] can be divided into two categories. The first category [125, 126, 128] focuses on creating augmented views through human design. Zhou et al. [125] introduce a Self-supervised Mobility Learning (SML) framework to encode human mobility semantics and enhance location-based tasks. SML addresses challenges posed by sparse and noisy mobility data by leveraging spatio-temporal contexts and augmented traces. SML generates contrastive views via traditional ways like sampling trajectory points. Another work [126] introduces a paradigm that contrastively learns trajectory representations from geolocation and temporal transitions via random sampling POI points to obtain contrastive views, enhancing the model’s understanding of mobility patterns. At its core is a Graph-enhanced Self-attentive layer, which uncovers transitional dependencies among POIs and captures temporal interests. Lastly, SelfTrip [128] focuses on trip recommendation using a two-step contrastive learning mechanism to learn nuanced representations of POIs and itineraries based on generated views via random walk, further enhanced by four innovative trip augmentation methods.
Nonetheless, recent studies [129, 44, 131, 132] on the self-supervised paradigm focus on creating augmented views automatically. While many existing studies prioritize final POI recommendation performance, they often overlook the crucial connection between POI sequences and contextual information. This can lead to limited data representations and less accurate recommendations. To address this, Jiang et al. [129] propose LSPSL, a unified attention framework for next POI recommendation that leverages self-supervised learning to model users’ long- and short-term preferences. LSPSL utilizes a self-attention network and two self-supervised optimization objectives to deeply explore the inherent relationships between POI sequences and contextual information and generate contrastive views automately during a pre-training phase. This process strengthens the data representations, allowing the model to better understand the nuances of user behavior. GSBPL [44] has been introduced for the domain of next-POI recommendation. GSBPL utilizes a pair of graph data augmentation techniques based on modeling user behaviour patterns via GNNs, to create augmented trajectory graphs, enabling the effective modeling of implicit behavioral patterns. To acquire high-order representations of trajectory graphs, GSBPL incorporates a graph preference representation encoder that utilizes geographical and social context. By applying contrastive learning, GSBPL captures implicit behavior patterns within these representations. To address the limitations of existing approaches that neglect the complex higher-order relationships between user-POI dynamics over time and fail to account for the varying importance of POIs, leading to an inability to effectively capture geographical influence imbalances, Fu et al. [132] introduce SLS-REC, a novel self-supervised long- and short-term model for POI recommendation. SLS-REC leverages a spatio-temporal attention hypergraph neural network to capture the spatial dependencies and temporal evolution of the short-term dynamic interests of users and creates contrastive views automately.
IV-D Generative Model-Based
IV-D1 Large Language Model-Based
Large language models are advanced deep learning models with millions of parameters, excelling at understanding and generating human language with high accuracy and fluency. They perform tasks like translation, summarization, and question-answering by predicting word sequences. Depending on their architectural focus, LLMs can be broadly categorized into three types: encoder-only [164], decoder-only [165], and encoder-decoder models [166]. Encoder-only LLMs focus on processing input text to create context-aware representations, excelling in tasks like classification, question answering, and information retrieval. Decoder-only LLMs specialize in text generation, predicting the next word in a sequence for tasks such as story writing and summarization. Encoder-decoder LLMs combine both components, processing input with an encoder and generating output with a decoder, making them ideal for tasks requiring both comprehension and generation.
Within the realm of POI recommendation, recent studies showcase the promising capabilities of LLMs. Feng et al. [46] explore the use of LLMs, specifically ChatGPT, for predicting a user’s next check-in location. Their work focuses on: (1) Prompt Engineering: Designing effective prompting strategies to elicit accurate location predictions from the LLM; (2) Ranking-Based Recommendation: Framing the recommendation task as a ranking problem, where the LLM ranks potential POIs based on their likelihood of being the user’s next destination; and (3) Incorporating Spatio-Temporal Factors: Considering key factors influencing human mobility, such as user preferences, spatial distances, and sequential transitions, in the prediction process. Different from [46], Li et al. [47] investigate how pretrained LLMs can effectively leverage the rich contextual information embedded within LBSN data. This method aims to: (1) Preserve Data Heterogeneity: Maintaining the original format of LBSN data, thereby preventing the loss of valuable contextual information often encountered during preprocessing steps; and (2) Context-Aware Representation Learning: Utilizing the LLM’s ability to capture complex relationships within unstructured data to learn more nuanced and context-aware representations of users and POIs.
IV-D2 Diffusion Model-Based
Diffusion models (DMs) [167] are a class of generative models that create data by gradually transforming a simple distribution into a complex one through a sequence of stochastic steps. Denoising Diffusion Probabilistic Models (DDPMs) [25], a specific type of DM, refine this process by learning to reverse the diffusion, progressively removing noise from a noisy signal to generate high-quality samples. DDPMs consist of two main phases: the forward and the reverse diffusion process, shown in Fig. 4 (b).
DMs have emerged as a powerful tool for generating high-quality, diverse data across various domains. Recent research explores their application to POI recommendation [133, 134, 135, 136], leveraging their ability to model complex distributions and capture intricate patterns in user behavior. Long et al. [136] introduce DCPR, a collaborative learning framework utilizing diffusion models for the next POI recommendation. Key features include a Cloud-Edge-Device architecture, region-specific recommendations, and collaborative learning. Zuo et al. [135] propose Diff-POI, a model leveraging diffusion models to capture users’ spatial preferences for POI recommendation. Notable aspects include spatial preference modeling, custom graph encoding, and diffusion-based sampling techniques. Wang et al. [134] present DSDRec, a recommender system combining diffusion models with semantic information extraction for next POI recommendation. Key highlights include the integration of semantic information extraction, pretrained language model enhancements, and spatio-temporal area encoding.
IV-D3 Variational Autoencoder
The variational autoencoder (VAE) [48, 49] is a unique type of autoencoder that introduces regularization to the encoding distribution during training, resulting in a well-structured latent space. VAEs offer a flexible framework for learning deep latent-variable models and their associated inference mechanisms.
Different VAEs [137, 138] offer unique advantages in POI recommendation, addressing various aspects of user behavior and data characteristics. WaPOIR [137] emphasizes understanding the data distribution to enhance personalization. It integrates diverse data sources, including user preferences, social influences, and geographical information, employing a Wasserstein autoencoder to learn latent features of users and POIs, effectively capturing the overall data distribution. The model captures both long-term preferences from historical check-ins and short-term inclinations from recent visits. Li et al. [138] employs a grid-based structure to incorporate spatial context, organizing check-in records and embedding users within their grid cells using a VAE. The decoder reconstructs check-in patterns, and a classifier associates representations with users, facilitating a structured approach to user modeling.
TABLE II: Model comparison in terms of effectiveness, efficiency, scalability and explainability
Criterion LFMs NNs SSL Gen-Models
Model Effectiveness Moderate High High High
Inference Efficiency High Varies Varies Low
Training Scalability High Moderate Moderate Low
Result Explainability Moderate Low Moderate High
IV-E Discussion of Models
In this part, we further compare the discussed model types: Latent Factor Models (LFMs), Classic Neural Networks (NNs), Self-Supervised Learning (SSL), and Generative Models (Gen-Models) from the aspects of model effectiveness, inference efficiency, training scalability, result explainability, shown in Table II. The comparison analysis is shown as:
Model effectiveness:
  • LFMs: LFMS moderately captures user-POI interactions using latent features for comparable recommendations.
  • NNs: NNs potentially achieve higher effectiveness by learning complex non-linear relationships between users and POIs, leading to more nuanced recommendations.
  • SSL: The effectiveness of SSL can be high, particularly when labeled data is scarce, by learning robust representations from unlabeled data.
  • Gen-Models: Gen-Models obtain high potential effectiveness due to their ability to generate diverse POI recommendations based on learned data distributions.
Inference efficiency:
  • LFMs: LFMS have high prediction efficiency, especially for smaller datasets, resulting in small prediction time.
  • NNs: Efficiency of NNs varies greatly depending on the network architecture and size; larger, more complex networks are less efficient.
  • SSL: Efficiency depends on the specific method employed; some approaches are more efficient than others.
  • Gen-Models: The inference efficiency of Gen-Models lower than other kinds of models due to their high model complexity and large model volume.
Training scalability:
  • LFMs: Scalability of training is higher than other models due to its simple model structure.
  • NNs: Training scalability is moderate; it’s limited only when the computational resources required for training complex networks become insufficient.
  • SSL: SSL often exhibits similar training scalability as NNs, especially methods using simpler architectures.
  • Gen-Models: Scalability is a major challenge; training scalability becomes computationally expensive with increasing data size.
Result explainability:
  • LFMs: Results of LFMs offer moderate explainability; latent factors provide some insights into user preferences.
  • NNs: Explainability is generally low; understanding results of complex neural networks is difficult.
  • SSL: Explainability is moderate; interpreting the learned representations requires further analysis.
  • Gen-Models: Explainability is higher than results of other models; understanding generated recommendation results often necessitates analyzing the latent space.
V Architecture
In this section, we provide illustrations on architectures, including centralized-based, decentralized-based, and federated learning-based architectures.
V-A Centralized Architecture
Centralized architecture serves as a robust, scalable, and cost-effective foundation for building and deploying recommendation systems. In this model, centralized data centers and servers provide a consolidated platform for managing large-scale computational resources. By centralizing these resources, enterprises can swiftly respond to changes in demand, reduce overhead, and streamline system maintenance. A key benefit of this architecture is its ability to centralize data processing and storage, allowing for more efficient management of data flow and integration across systems. The flexibility of centralized architecture can be expressed through the following equation:
[…]
Scalability+Cost Efficiency+AccessibilityInitial Hardware Investments+Physical $\frac{\text{Scalability}+\text{Cost Efficiency}+\text{Accessibility}}{\text{Initial Hardware Investments}+\text{Physical Constraints}}$
This equation highlights how centralized systems optimize scalability, cost-effectiveness, and resource availability, enabling rapid adaptation without requiring extensive upfront hardware investments. We explore studies [51, 50, 88, 139, 53] on POI recommendation systems developed using centralized architecture. For instance, Wang et al. [51] leverage centralized computing to model geographical influences in POI recommendations with GeoIE. Centralized servers enable real-time processing of geographical data, enhancing the efficiency of recommendations by factoring in geo-influence, geo-susceptibility, and physical distance. Similarly, Yin et al. [50] developed SH-CDL, where centralized architecture plays a crucial role in managing the fusion of diverse POI features and user preferences across multiple locations. The centralization of data allows SH-CDL to more effectively model both individual and public preferences by leveraging collective data from a large number of users. Furthermore, Wang et al. [53] introduced AGRAN, which benefits from centralized data integration to create adaptive POI graphs and fuse POI representations with dynamic spatial-temporal data, ensuring real-time responsiveness to evolving user preferences. These studies underscore the advantages of centralized architecture in POI recommendation systems, emphasizing enhanced efficiency and scalability. However, as the complexity of these systems increases, continuous innovation in defense mechanisms is essential to safeguard data integrity and user privacy within centralized frameworks.
V-B Decentralized-based Architecture
The architecture of decentralized methodologies prioritizes the maximal utilization of local storage and processing capabilities, thereby conserving bandwidth and curtailing cloud storage expenses. Moreover, by processing data locally, these methodologies bolster data security and user privacy, mitigating the risks linked to data breaches and unauthorized access. Mathematically, the efficacy of decentralized methods can be encapsulated in the equation:
[…]
Local Processing Power+Storage CapacityBandwidth Consumption+Cloud Storage $\frac{\text{Local Processing Power}+\text{Storage Capacity}}{\text{Bandwidth Consumption}+\text{Cloud Storage Costs}}$
where the decentralized efficiency metric integrates local processing power, storage capacity, bandwidth consumption, and cloud storage expenses to gauge the overall effectiveness of decentralized computing solutions in comparison to cloud-based alternatives.
TABLE III: Related works in different application sub-tasks. We provide detailed information on the names of journals/conferences, the specific generative techniques used, the particular application sub-tasks addressed, and the datasets used for evaluation.
Models Source Technique Sub-task Datasets
Latent Factor Model-based Models
LDA-based Models
LCLCARS [64] KDD 2013 LDA Location-Content-Aware Recommender Foursquare, and Douban Event
LCARS [16] TOIS 2014 LDA Spatial Item Recommender Douban Event, and Foursquare
JIM [65] CIKM 2015 LDA POI Recommendation Foursquare, and Twitter
Geo-SAGE [66] KDD 2015 LDA Spatial Item Recommendation Foursquare, and Twitter
TRM [65] ACM MM 2015 LDA POI Recommendation Foursquare, and Twitter
LA-LDA [67] TKDD 2015 LDA Location-based Recommendation MovieLens, Gowalla, and Douban Event
TRM+ [68] TOIS 2016 LDA POI Recommendation Foursquare, and Twitter
SPORE [69] ICDE 2016 LDA Spatial Item Recommendation Foursquare, Twitter [168], and Synthetic Dataset
ST-LDA [70] TKDE 2016 LDA POI Recommendation Foursquare, and Gowalla
UCGT [71] ICDE 2016 LDA Social Community Detection Foursquare, and Douban Event
LSARS [72] KDD 2017 LDA Spatial Item Recommendation Yelp, and Foursquare
ST-SAGE [73] TIST 2017 LDA Spatial Item Recommendation Foursquare, and Twitter
TPM [74] TIST 2018 LDA Spatial Item Recommendation Foursquare, Twitter [168], and Synthetic Dataset
MF-based Models
GeoMF [75] KDD 2014 MF POI Recommendation Jiepang
GeoMF++ [76] TOIS 2018 MF POI Recommendation Gowalla, and Jiapang
LGLMF [77] AIRS 2019 MF POI Recommendation Foursquare, and Gowalla
SSTPMF [78] KIS 2021 MF POI Recommendation Foursquare-TKY, and Gowalla-NYC
SQPMF [79] Applied Intelligence 2024 MF POI Recommendation GowallaFootnote, Foursquare, and Brightkite
Classic Neural Network (NN)-based Models
LSTM-based Models
ST-LSTM [52] CoRR 2018 LSTM POI Recommendation CA, Gowalla, SIN, Brightkite
LSTM-S [81] MDM 2019 LSTM POI Recommendation Foursquare
PS-LSTM [82] CNIT 2021 LSTM POI Recommendation Yelp, Gowalla, Foursquare
STMLA [83] ICWS 2022 LSTM POI Recommendation Gowalla, Foursquare
LSMA [84] IJGI 2022 LSTM Next POI Recommendation Foursquare-Charlotte (CHA) [169], Foursquare-NYC
AT-LSTM [85] IJSWIS 2023 LSTM POI Recommendation Gowalla, Foursquare
ATSD-GRU [86] IJITSA 2023 LSTM POI Recommendation mafengwo
DLAN [87] IFS 2024 LSTM POI Recommendation Foursquare-NYC, Foursquare-TKY
Transformer-based Methods
TLR-M [41] PAKDD 2021 Transformer POI Recommendation Foursquare-NYC, Foursquare-TKY
GETNext [88] SIGIR 2022 Transformer POI Recommendation Foursquare-NYC, Foursquare-TKY, CA
CARAN [89] IEEE Access 2022 Transformer Next POI Recommendation Foursquare-NYC, Foursquare-TKY, and Gowalla
JANICP [90] DSE 2022 Transformer POI Recommendation Weeplaces, Foursquare-NYC and Foursquare-TKY
Li et al. [91] IJGI 2022 Transformer Next POI Recommendation Foursquare-NYC and Foursquare-TKY
STUIC-SAN [92] IJGI 2022 Transformer Next POI Recommendation Foursquare, Gowalla
AMACF [93] IJGI 2022 Transformer Next POI Recommendation Foursquare‐NYC, Foursquare-TKY and Weeplaces
CHA [94] TOIS 2022 Transformer Next POI Recommendation Foursquare‐NYC, Foursquare-TKY
HAT [96] IEEE MM 2023 Transformer POI Recommendation BJ, SH, NJ, CD
STAR-HiT [97] ACM TIS 2023 Transformer POI Recommendation Foursquare‐NYC, Foursquare‐US, Gowalla
CAFPR [98] Applied SC 2023 Transformer POI Recommendation Foursquare-TKY, CA, Budapest [170], Melbourne [170], Magic k [170]
FPG [99] Applied SC 2023 Transformer POI Recommendation Foursquare, Gowalla
TGAT [100] IF Systems 2023 Transformer POI Recommendation Foursquare-NYC, Foursquare-TKY
MobGT [101] SigSpatial 2023 Transformer POI Recommendation Foursquare-NYC, Gowalla
POIBERT [102] IEEE Big data 2022 Transformer POI Recommendation Budapest, Delhi, Edinburgh, Glasgow, Osaka, Perth, Toronto
AutoMTN [103] SIGIR 2022 Transformer POI Recommendation Foursquare-NYC, Foursquare-TKY
CCDSA [104] Applied Intelligence 2023 Transformer Next POI Recommendation Foursquare-NY, Foursquare-TKY, Weeplaces-NY and Weeplaces-SF
Liu et al. [105] IJITSA 2023 Transformer POI Recommendation BrightKite
TDGCN [106] IPM 2023 Transformer Next POI Recommendation Foursquare-TKY, Weeplaces and Gowalla-CA
TGAT [100] JIFS 2023 Transformer POI Recommendation Foursquare-NYC, Foursquare-TKY
MGAN [107] JIFS 2023 Transformer POI Recommendation Yelp, Foursquare
STA-TCN [108] TKDE 2023 Transformer Next POI Recommendation Gowalla, Foursquare
Xia et al. [109] PerCom Workshops 2023 Transformer Next POI Recommendation Gowalla, Foursquare
STTF-Recommender [110] IJGI 2023 Transformer Next Location Recommendation Gowalla, Foursquare
BayMAN [111] TKDE 2023 Transformer POI Recommendation Gowalla, NYC, Foursquare
Kumar et al. [112] IF 2024 Transformer POI Recommendation Gowalla, Foursquare
ImNext [113] KBS 2024 Transformer Next POI Recommendation Gowalla, Foursquare
BSA-ST-Rec [114] MTA 2024 Transformer POI Recommendation Gowalla, Foursquare
Wu et al. [115] TMM 2024 Transformer POI Recommendation Ctrip, Baidu Travel, Qunar, Yelp
ROTAN [116] KDD 2024 Transformer Next POI Recommendation Foursquare-NYC, Foursquare-TKY, Gowalla-CA
Graph Neural Network (GNN)-based Methods
STGCN [117] ICDM 2020 GCN POI Recommendation Yelp, Boston, Chicago, London
Dynaposgnn [20] ICDM Workshops 2021 GNN POI Recommendation Gowalla, Foursquare
ASGNN [21] World Wide Web 2021 GNN Next POI Recommendation Gowalla, FourSquare, and Brightkite
GNN-POI [118] Neurocomputing 2021 GNN POI Recommendation Gowalla, Foursquare, and Yelp
ADQ-GNN [22] WISE 2021 GNN POI Recommendation Foursquare-NYC, Foursquare-TKY, Gowalla
GN-GCN [119] ICIIW 2022 GCN POI Recommendation Gowalla, Yelp
ATST-GGNN [120] IEEE Access 2022 GNN Next POI Recommendation Foursquare, Gowalla
Zhao et al. [121] MDM 2023 GNN POI Recommendation Wi-fi login
HS-GAT [122] Expert Systems 2024 GNN POI Recommendation Yelp, Boston, Chicago, London
PROG-HGNN [28] Expert Systems 2024 GNN POI Recommendation Foursquare, Gowallam, and Yelp
CGNN-PRRG [123] IPM 2024 GNN POI Recommendation Foursquare, Gowalla, and Yelp
HKGNN [124] World Wide Web 2024 GNN Next POI Recommendation Foursquare-NYC, Foursquare-JK, Foursquare-KL, and Foursquare-SP
Self Supervised Learning (SSL)-based Models
SML [125] KBS 2021 SSL Next Location Prediction Foursquare, and Gowalla
S2GRec [126] arXiv 2022 SSL Next POI Recommendation Gowalla, Foursquare-TKY and Foursquare-NYC
SSTGL [127] Applied Sciences 2023 SSL POI Recommendation Foursquare, Gowalla, and Meituan
SelfTrip [128] KBS 2022 SSL Trip Recommendation Toronto, Osaka, Glasgow, and Edinburgh
GSBPL [44] Electronics 2023 SSL Next POI Recommendation Gowalla, Foursquare-TKY, and Foursquare-NYC
LSPSL [129] TKDD 2023 SSL Next POI Recommendation Foursquare-NYC and Foursquare-TKY
Gao et al. [130] TKDE 2023 SSL Human Trajectory Prediction Foursquare-NYC, Foursquare-TKY, Los Angeles, Houston
Wang et al. [131] CSCWD 2024 SSL POI Recommendation Foursquare, and Gowalla
SLS-REC [132] Expert Systems 2024 SSL POI Recommendation Foursquare, and Gowalla
CLLP [45] SIGIR 2024 SSL Next POI Recommendation Foursquare, and Gowalla
Generative-based Methods
LLM-based Methods
LLMmove [46] Arxiv 2024 LLM POI Recommendation NYC [171], TKY [171]
Li et al. [47] SIGIR 2024 LLM POI Recommendation NYC [171], TKY [171], Gowalla-CA
Diffusion Models (DMs)-based Methods
Diff-POI [133] TOIS 2023 Diffusion Next POI Recommendation Gowalla, Foursquare-SIN, Foursquare-TKY, and Foursquare-NYC
DSDRec [134] Information Sciences 2024 Diffusion Next POI Recommendation Foursquare-NYC, and Foursquare-TKY
Diff-DGMN [135] IOT 2024 Diffusion POI Recommendation IST, JK, SP, NYC, LA
DCPR [136] arXiv 2024 Diffusion POI Recommendation Foursquare, and Weeplace
VAE-based Methods
WaPOIR [137] TSMCS 2023 VAE POI Recommendation Gowalla, Foursquare
Li et al. [138] ICAI 2022 VAE POI Recommendation Gowalla, Foursquare
In this part, we illustrate the decentralized architecture for POI recommendation. While centralized architectures have excelled in accuracy, they require substantial computational resources during the training phase. To address these challenges, researchers have introduced decentralized architectures like [54, 27, 140]. To be specific, Long et al. [54] propose an innovative method for decentralized POI recommendations called Model-Agnostic Collaborative Learning for decentralized POI recommendation, which empowers users to tailor their own model architectures, such as adjusting dimensions and the quantity of hidden layers. To address the scarcity of user data on devices, MAC [54] introduces a strategy to pre-determine collaborators based on physical proximity, preferences at the category level, and social connections. For efficiently and securely integrating insights from these chosen collaborators, MAC implements a knowledge distillation approach with a focus on maximizing mutual information. In MAC, rather than exchanging sensitive models or gradients, clients selectively share their informed decisions concerning a predetermined reference dataset.
However, current collaborative learning systems commonly rely on a single reference dataset shared among all users, which may not cater to individual preferences, potentially hindering knowledge exchange and overall performance. Users may struggle to provide relevant insights on items outside their scope of interest, further complicating the collaborative process. To bridge this gap, Zheng et al. [140] propose the Adaptive Reference Data for Decentralized Collaborative Learning framework. DARD tailors adaptive reference datasets for enhanced user collaboration by creating a desensitized public reference pool using transformation and probability data generation techniques. For each user, DARD dynamically selects personalized reference data through loss tracking and influence function analysis. Individual models are trained using private data and collaboratively with spatial and contextual neighbors. During collaboration, users exchange soft decisions based on a blend of their personalized adaptive reference data, fostering more effective collaboration and personalized recommendations.
V-C Federated Learning-based Architecture
Federated Learning (FL) [55] revolutionizes machine learning model training by enabling a decentralized and collaborative approach that ensures data privacy. This innovative technique allows models to be trained locally on devices without sharing raw data, thus providing a secure solution for deploying new machine learning applications. Fig. 6 presents a comparison between the centralized, decentralized, and FL-based architectures.
The FL-based approach offers the advantage of reducing data transfer and mitigating privacy risks by keeping sensitive information on local devices. The efficacy of Federated Learning can be mathematically captured through the following equation:
[…]
DPP+Computational Resource OptimizationBandwidth Utilization+Model Update $\frac{\text{DPP}+\text{Computational Resource Optimization}}{\text{Bandwidth Utilization}+\text{Model Update Aggregation}}$
Where DPP denotes Data Privacy Protection. This formula encapsulates how Federated Learning optimizes data privacy protection, computational resources, bandwidth usage, and model update aggregation to bolster the efficiency of machine learning model training across distributed devices.
Figure 5: Data poisoning (Left) and model poisoning (Right)
FL operates on the principle of training machine learning models without requiring access to individual users’ raw private data [142, 143, 144, 145, 146, 29, 147, 148, 149, 150]. An earlier pioneering work [142] introduces a cross-domain POI framework that incorporates federated learning techniques and privacy safeguards. By leveraging data from a related domain in user interest profiling, it aims to mitigate the cold-start issue. Additionally, the framework employs federated learning to locally analyze users’ historical data and encrypts latent feature distributions to ensure user privacy while transferring knowledge. The method aims to solve challenges like the sensitive and sparse nature of POI data and the cold-start dilemma stemming from data scarcity, which hampers recommendation effectiveness.
Another earlier work [148] introduces PREFER, a novel federated learning framework optimized for POI recommendation with edge acceleration. PREFER divides the recommendation process into distinct segments. Initially, for privacy preservation, users develop individualized recommendation models locally and exchange multi-dimensional user-agnostic parameters instead of sharing check-in data. Subsequently, to enhance recommendation speed, these decentralized parameters are consolidated on edge servers situated near users (like base stations), as opposed to distant cloud servers. Besides, Huang et al. [143] present a novel geographical Point of Interest recommendation approach based on federated learning. The POI recommendation task is framed as a matrix factorization optimization challenge, with the application of singular value decomposition for matrix decomposition. By demonstrating the non-convex nature of the optimization problem, the framework integrates stochastic gradient descent within SVD and devises a FL framework for concurrently addressing the POI recommendation task. Within our FL strategy, solely the computed gradient details are transmitted from users to the FL server, while users independently handle their rating and geographic preference data on their personal devices to ensure privacy during communication.
FedPOIRec [29] is a privacy-focused framework that ensures local data remains on the user’s device while updates are aggregated blindly by a parameter server. To enhance personalization, FedPOIRec enables users to exchange learned parameters with their friends, facilitating knowledge transfer. To further protect privacy, we introduce a protocol that integrates the preferences of a user’s friends after federated computation. This protocol leverages the properties of the CKKS fully homomorphic encryption scheme to ensure privacy during this integration process.
Figure 6: The centralized, decentralized, and FL-based approaches.
VI Security
In this section, we provide illustrations on security, including data integrity threats and user privacy protection in POI recommender systems.
VI-A Data Integrity Threats
Data integrity threats refer to malicious activities that compromise the accuracy, consistency, and trustworthiness of data within systems, particularly those that rely heavily on data-driven technologies. These threats can manifest at various stages, including data storage, transmission, and processing, significantly impacting the utility of POI recommender systems.
One prominent data integrity threat in POI recommendation systems is poisoning attacks [56, 57]. These attacks manipulate training data, undermining the credibility and effectiveness of machine learning systems by injecting malicious data. Such attacks skew model predictions to benefit the attacker. Figure 5 illustrates the mechanisms of data and model poisoning, highlighting vulnerabilities from compromised datasets. The influence of poisoning attacks can be formally represented by manipulating training data, altering the learning process of machine learning models, as shown in the equation $\mathbf{X}_{\mathsf{poisoned}}=\mathbf{X}_{\mathsf{clean}}+\mathbf{A}_{\mathsf{perturb}}$, where $\mathbf{X}_{\mathsf{poisoned}}$ denotes the poisoned data, $\mathbf{X}_{\mathsf{clean}}$ represents the clean data, and $\mathbf{A}_{\mathsf{perturb}}$ signifies the adversarial perturbation introduced during the attack.
Rong et al. [56] introduced FedRecAttack, a model poisoning attack targeting federated recommendation systems. By exploiting public interactions to approximate user feature vectors, attackers upload poisoned gradients through malicious users. Experiments show FedRecAttack is highly effective, even with minimal malicious users and public interactions, revealing FR’s unexpected vulnerability. Zhang et al. [57] presented LOKI, a data poisoning attack on next-item POI recommendation systems. Using reinforcement learning, LOKI generates adversarial user behaviors, reducing retraining costs by interacting with a simulator. Results show LOKI surpasses previous attacks across diverse POI datasets and models. Another study et al. [151] introduced a new poisoning attack method that manipulates the ranking and exposure of top-K recommended targets in recommendation systems. Unlike previous methods, this attack operates without prior system knowledge, employing synthetic malicious users who strategically upload poisoned gradients targeting products related to desired items.
Additionally, POI recommendation systems may face various other threats, including Sybil attacks and the physical trajectory inference attack. Wang et al. [58] revealed vulnerabilities in real-time crowdsourced maps, demonstrating how software-based Sybil devices can exploit weak location authentication. Their techniques enable large-scale generation of these devices, leading to significant security and privacy threats in systems like Waze. Recently, Long et al. [59] introduced a novel physical trajectory inference attack (PTIA) that exploits aggregated information from POIs to reveal users’ historical trajectories in decentralized recommendation.
These studies highlight the increasing sophistication of attacks on POI recommendation systems, revealing critical vulnerabilities. As attackers refine their methods, robust and adaptive defenses must evolve to protect the integrity and trustworthiness of POI recommendation processes.
VI-B User Privacy Protection
In recent years, privacy concerns have become critical issues in POI recommendation systems. These systems systems process vast amounts of sensitive user information to generate personalized recommendations, making them attractive targets for malicious actors seeking to exploit this data.
To address privacy concerns, various cryptographic techniques and privacy-preserving mechanisms are essential for safeguarding sensitive information in POI recommendation systems. Decentralized secure computation [60, 61] enables multiple parties to collaboratively compute a function over their inputs while keeping them private. Chen et al. [60] propose a decentralized matrix factorization framework for POI recommendation, addressing privacy and computational issues by enabling decentralized training on user devices. This approach preserves users’ rating data and enhances recommendation accuracy. Furthermore, Chen et al. [61] introduce the PriRec framework, which maintains users’ private data on their devices while employing local differential privacy techniques and secure decentralized gradient descent to protect model privacy, achieving comparable recommendation accuracy.
Figure 7: The workflow of homomorphic encryption
Additionally, homomorphic encryption [63, 29] is also commonly used for privacy-preserving POI recommendation. It allows computations on encrypted data without decryption, ensuring sensitive information remains secure. Figure 7 illustrates the workflow of homomorphic encryption, which conceals the underlying plaintext during processing. Li et al. [63] develop a privacy-preserving framework for POI recommendation by utilizing homomorphic encryption. They address the cold start problem prevalent in POI recommendation systems, which arises due to the sparsity of user-location check-in data. Their framework leverages partially homomorphic encryption to enable trust-oriented recommendations while ensuring that sensitive user check-in and trust data remain private. By employing offline encryption and parallel computing, Li et al. [63] ensure that their protocols efficiently protect the private data of all parties involved in the recommendation process. They also prove that their proposed protocols are secure against semi-honest adversaries, and their experiments on both synthetic and real datasets demonstrate the feasibility of privacy-preserving recommendations with acceptable computation and communication costs.
VII Future Directions
As POI recommendation systems continue to evolve, vast opportunities for innovation will keep emerging across the three pivotal dimensions: models, architectures, and security. By integrating cutting-edge model structures, adopting flexible architectures, and reinforcing security measures, the future of POI recommendation systems are expected to be more adaptive, personalized, and resilient to emerging challenges. Below, we outline the key future directions that will shape the development of these systems across these essential aspects.
Future POI recommendation models will move beyond traditional and deep learning techniques, incorporating various data modalities and more advanced methodologies to further improve accuracy and personalization. (1) LLM Agent-Powered POI Recommendation. The integration of Large Language Model (LLM) agents [172] into POI recommendation systems offers another approach to enhance personalization POI recommendation. LLM agents can interpret user preferences through natural language interactions, making real-time POI recommendations based on conversational queries and feedback [173, 174]. These agents can provide personalized POI suggestions by engaging users in dialogue to clarify needs and generate tailored recommendations. By incorporating contextual factors such as location, time, and user-specific constraints, LLM-powered POI recommendation systems offer dynamic, context-aware suggestions. (2) LLM-Powered Explainable POI Recommendation. Creating an explainable POI recommendation system via large language models is a promising avenue for leveraging advanced language processing capabilities to provide transparent and understandable POI recommendations. By utilizing these models, users can receive recommendations accompanied by explanations that clarify why a particular POI is suggested. This transparency enhances user trust and engagement by shedding light on the decision-making process behind each recommendation, ultimately improving the overall user experience of POI recommendation.
Future POI recommendation architecture is to ensure that decentralized systems can deliver the same level of personalization and accuracy as their centralized counterparts, all while maintaining the privacy benefits that decentralized architectures provide. Small Language Model-Powered Decentralized POI Recommendation. While the mainstream focus is on developing large language models, small language models are also gaining attention for their efficiency and ability to run on-device, offering privacy-preserving and real-time capabilities. The newly released Llama 3.2’s lightweight, language models (1B and 3B)  [^1] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ can revolutionize edge AI and significantly improve the accuracy of decentralized POI recommendation systems by processing real-time, contextual data directly on the device. These models can learn from user interactions, preferences, and location data locally by using the long contexts (up to 128K tokens), allowing for highly personalized recommendations without needing centralized data aggregation. By running on-device, the models can continually adapt to user behavior, leading to more precise recommendations based on immediate context, while preserving privacy. The emergence of new models and architectures has brought up new opportunities, consequently expanding the landscape of potential threats. Future POI recommendation systems need to handle new security issues. (1) Addressing Vulnerabilities in Emerging POI Recommender Systems. As novel techniques such as multi-modal data infusion, self-supervised learning, and LLM-based recommendation systems gain traction, they introduce new layers of complexity and new vulnerabilities. The security landscape of POI recommendation
systems will need to focus on understanding these emerging attack surfaces, as they differ significantly from traditional techniques. Future research should prioritize identifying and mitigating vulnerabilities unique to these advanced techniques, such as adversarial manipulation in multi-modal inputs or exploiting context-driven weaknesses in LLM-based systems. By developing robust security frameworks tailored to these innovative approaches, future POI recommendation systems can safeguard against both known and unforeseen threats, ensuring trust and reliability as these technologies evolve. (2) Privacy-Enhancing POI Recommendation based on Cryptography Techniques. Developing privacy-enhancing techniques for POI recommendation systems is crucial for protecting user data while maintaining recommendation quality. This involves implementing methods such as secure multiparty computation and zero-knowledge proof, ensuring user interactions and preferences remain confidential. By safeguarding sensitive information, these techniques can enhance user trust and encourage broader adoption of recommendation systems, thereby balancing privacy protection with system performance and ensuring that privacy-enhanced models do not compromise the accuracy and relevance of recommendations. By embracing these diverse research directions and leveraging cutting-edge technologies and methodologies, the field of POI recommendation is poised to revolutionize the way users discover and engage with points of interest in the digital landscape.
VIII Conclusion
The abundance of spatio-temporal data generated by smartphones and location-based social networks has fueled advancements in POI recommendation systems. These systems enhance user experiences, personalize interactions, and streamline decision-making in the digital realm. Our survey bridges the gap in existing literature by providing a comprehensive overview of recent advancements in POI recommendation, encompassing models, architectures, and security aspects. We highlight the emergence of advanced machine learning models, explore evolving system architectures, and emphasize the crucial need for robust security measures to protect user privacy and prevent malicious manipulation. This comprehensive taxonomy serves as a valuable resource for researchers and practitioners seeking to advance the field of POI recommendation and harness the full potential of spatio-temporal data.

### v2

@@
 The proliferation of smart devices has fueled the rapid growth of Location-based Social Networks (LBSNs) [1, 2, 3], which allow users to share check-ins, reviews, and personal experiences tied to specific locations.
-These networks, now with billions of users, generate vast amounts of spatio-temporal data [4, 5, 6, 7, 8, 9], presenting valuable opportunities for personalized Point-of-Interest (POI) recommendations.
+These networks, now with billions of users, generate vast amounts of spatio-temporal data [4, 5, 6, 7, 8, 9, 10, 10, 11], presenting valuable opportunities for personalized Point-of-Interest (POI) recommendations.
 As a dynamic area in recommendation systems, POI recommendation has gained considerable interest from both users and businesses in recent years.
@@
 These techniques have shifted the landscape towards more robust systems capable of balancing both security and performance.
-While existing surveys [34, 35, 36, 37] have contributed valuable insights into POI recommendation systems, there remains a critical need for a comprehensive review that reflects the rapid developments across POI models, architectures, and security.
+While existing surveys [35, 36, 37, 38, 39] have contributed valuable insights into POI recommendation systems, there remains a critical need for a comprehensive review that reflects the rapid developments across POI models, architectures, and security.
 For instance, while Zhao et al.
-[37] provide a thorough review of POI recommendation with traditional techniques like matrix factorization, the growing challenges within deep POI recommender systems are not covered.
+[35] provided a thorough review of POI recommendation with traditional techniques like matrix factorization, the growing challenges within deep POI recommender systems are not covered.
 By contrast, Wang et al.
-[34] provide an overview of various POI recommendation methods in the deep learning era, yet they do not delve deeply into the architectural and security challenges that have emerged with the rise of decentralized systems and practical privacy concerns.
+[38] provided an overview of various POI recommendation methods in the deep learning era, yet they do not delve deeply into the architectural and security challenges that have emerged with the rise of decentralized systems and practical privacy concerns.
 Similarly, Islam’s survey [37] emphasizes the impact of deep learning on POI recommendation, but it overlooks key advancements in federated learning and edge computing, which are increasingly shaping the deployment of these systems.
 Furthermore, Werneck’s survey [36] provides a detailed account of POI recommendation techniques from 2017 to 2019, offering valuable insights into the evolution of methodologies during that time frame, but it lacks coverage of more recent advancements such as the integration of GNNs and LLMs, which bring about not only powerful user modeling capabilities but also intensive computation and scalability issues.
-To summarize, a noticeable gap in the literature regarding cutting-edge models, architectural evolution, and security considerations is shown in Table I.
-TABLE I: In comparing our survey to related surveys in the field, several key distinctions and contributions emerge.
-Survey [37] [36] [35] [34] Ours
-Taxonomy No Yes Yes Yes Yes
-Models Yes Yes Yes Yes Yes
-Architecture No No No No Yes
-Security No No No No Yes
-Year 2016 2020 2022 2023 2024
+Recently, Yin et al.
+[39] provided a comprehensive survey on on-device recommender systems, covering aspects such as deployment, inference, training, updating, and the security and privacy of these systems.
+Consequently, their work does not address key aspects of POI recommendation, such as the foundational models used.
+As shown in Table I, there is a noticeable gap in the literature regarding cutting-edge models, architectural evolution, and security considerations.
+TABLE I: Comparison of related surveys.
+Survey Year Taxonomy Models Architecture Security
+[35] 2016 $\times$ $\checkmark$ $\times$ $\times$
+[36] 2020 $\checkmark$ $\checkmark$ $\times$ $\times$
+[37] 2022 $\checkmark$ $\checkmark$ $\times$ $\times$
+[38] 2023 $\checkmark$ $\checkmark$ $\times$ $\times$
+[39] 2024 $\checkmark$ $\times$ $\checkmark$ $\checkmark$
+Ours 2024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 With the increasing diversity of data sources, the rise of new models and architectures, and the mounting need for privacy-preserving techniques, it is essential to provide an up-to-date, comprehensive survey that covers these critical aspects.
@@
 The primary objective is to provide personalized recommendations of relevant POIs based on users’ past interactions and preferences within the LBSN.
-The key concepts of POI recommendation are articulated as follows:
+Table II provides a concise description of the primary notations used in this paper.
+TABLE II: Notations used throughout this paper.
+Notation Description
+$u_{i}$ User $i$ where $i\in\{1,2,...,N\}$
+$U$ Set of $N$ users within the LBSN, where $U=\left\{u_{1},u_{2},\ldots,u_{N}\right\}$.
+$p_{i}$ POI, a specific location recognized for its significance or interest.
+$P$ Set of $M$ POIs, where $P=\{p_{1},p_{2},...,p_{N}\}$.
+$\tilde{U}$ Network of user connections within the LBSN, where $\tilde{U}=\left\{(u_{i},u_{j})|u_{i},u_{j}\in U\right\}$.
+$x_{p}$ Latitude of POI $p$.
+$y_{p}$ Longitude of POI $p$.
+$A_{p}$ Set of attributes describing the semantics of POI $p$.
+$c^{u}_{o}$ Check-in event where user $u$ declares presence at a location at time $o$.
+$C^{u}$ List of check-in events for user $u$, where $C^{u}=\left\{c^{u}_{1},c^{u}_{2},\dots,c^{u}_{k}\right\}$.
+The key concepts of POI recommendation are as follows:
   • Point-of-Interest: A point of interest, represented as $p_{i}$, signifies a particular location or site recognized for its significance or interest to various individuals, tourists, or researchers.
 These points on a map hold specific value or importance and can encompass landmarks, attractions, businesses, historical sites, parks, restaurants, hotels, or any locale of potential interest.
-  • Check-in Event: A check-in event, denoted as $c^{u}_{o}$, involves a user publicly declaring their presence at a specific venue or location at a specific time step $o$.
-Typically facilitated through mobile applications or social media platforms, this action entails sharing the user’s current whereabouts with their social circle.
-During a check-in, the user typically selects or searches for a particular place, such as a restaurant, park, or event space, and indicates their current presence at that spot.
+  • Check-in Event: A check-in event, denoted as $c^{u}_{o}$, involves a user publicly declaring their presence at a venue or location at a specific time step $o$.
+Typically facilitated through mobile apps or social media, this action shares the user’s whereabouts with their social circle.
+During a check-in, the user selects or searches for a place, such as a restaurant, park, or event space, to indicate their presence.
+  • POI Recommendation: Given a user $u$’s check-in list $C^{u}=\left\{c^{u}_{1},c^{u}_{2},\dots,c^{u}_{k}\right\}$, where each $c^{u}_{i}$ represents a check-in event at time $o_{i}$, the POI recommendation task aims to suggest a set of POIs to the user based on their historical check-ins and preferences.
+The goal is to suggest POIs that the user is likely to visit in the future, leveraging their interaction patterns and preferences within the LBSN.
   • Next POI Recommendation: Given a user $u$’s check-in list $C^{u}$, the next POI recommendation task revolves around predicting the subsequent POI at time $o_{i+1}$.
+It specifically refers to recommending the immediate next POI the user is likely to visit based on their current trajectory and patterns.
+  • Spatial Item Recommendation: Given a user’s historical interaction data at various spatial locations, the spatial item recommendation task aims to suggest items or products that are contextually relevant to the user based on their location and spatial preferences.
+The goal is to recommend items that the user is likely to engage with, considering both the spatial distribution of items and the user’s past interactions within these geographical regions.
+{forest}
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=center, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [POI Recommendation,leaf-head, ver [ Models (Sec.
+IV), leaf-task,text width=8em [ Latent Factor Model-Based, leaf-task10, text width=11.5em [LDA-Based, leaf-task8, text width=7.5em[LCARS [40, 17], JIM [41], Geo-SAGE [42], TRM [41],
+LA-LDA [43], TRM+ [44], SPORE [45], ST-LDA [46], UCGT [47],
+LSARS [48], ST-SAGE [49], TPM [50], modelnode-task8,text width=29em]] [MF-Based, leaf-task9, text width=7.5em[GeoMF [51], GeoMF++ [52], LGLMF [53], SSTPMF [54],
+SQPMF [55], modelnode-task9,text width=29em]] ] [ Classic NN-Based, leaf-taska, text width=9.5em [LSTM-Based, leaf-task1, text width=9.5em [ST-LSTM [56], LSTM-S [57], PS-LSTM [58], STMLA [59],
+LSMA [60], AT-LSTM [61], ATSD-GRU [62], DLAN [63],modelnode-task1, text width=29em] ] [Transformer-Based, leaf-task2, text width=9.5em [TLR-M [64], GETNext [65], CARAN [66], JANICP [67],
+Li et al. [68], STUIC-SAN [69], AMACF [70], CHA [71],
+MVSAN [72], HAT [73], STAR-HiT [74], CAFPR [75],
+FPG [76], TGAT [77], MobGT [78], POIBERT [79],
+AutoMTN [80], CCDSA [81], Liu  et al. [82], TDGCN [83],
+TGAT [77], MGAN [84] , STA-TCN [85] , Xia et al. [86],
+STTF-Recommender [87] , BayMAN [88], Kumar et al. [89],
+ImNext [90] , BSA-ST-Rec [91] , Wu et al.  [92] , ROTAN [93],modelnode-task2, text width=29em] ] [GNN-Based, leaf-task3, text width=9.5em [STGCN [94], Dynaposgnn [21] , ASGNN [22] , GNN-POI [95] ,
+ADQ-GNN [23] , GN-GCN [96] , ATST-GGNN [97] , Zhao et al.
+ [98] , HS-GAT [99] , PROG-HGNN [29] , CGNN-PRRG [100] ,
+HKGNN [101],modelnode-task3, text width=29em] ] ] [ SSL-Based Models, leaf-task6, text width=9.5em [SML [102], S2GRec [103], SSTGL [104], SelfTrip [105], GSBPL [106], LSPSL [107],
+Gao et al.
+[108], Wang et al.
+[109], SLS-REC [110], CLLP [111], , modelnode-task6, text width=40em] ] [ Generative, leaf-taska, text width=9.5em [ LLM-Based, leaf-task5, text width=9.5em [LLMmove [112], Li et al.
+[113], modelnode-task5,text width=29em] ] [ DM-Based, leaf-task7, text width=9.5em [Diff-POI [114], DSDRec [115] , Diff-DGMM [116] , DCPR [117], modelnode-task7,text width=29em] ] [VAE-Based, leaf-task4, text width=9.5em [WaPOIR [118], Li et al.
+[119], modelnode-task4, text width=29em] ] ] ] [ Architectures
+(Sec.
+V), leaf-paradigms,text width=8em [ Centralized, leaf-paradigms, text width=9.5em [ GeoIE [120], SH-CDL [121], GETnext [65], STGN [122], AGRAN [123] , modelnode-paradigms, text width=40em] ] [ Decentralized, leaf-paradigms, text width=9.5em [ MAC [124], DARD [125], LLRec [126], DCLR [28], modelnode-paradigms, text width=40em] ] [ FL-Based, leaf-paradigms, text width=9.5em [ FL&PP-POI [127], Huang et al.
+[128],CPF-POI [129], SCFL [130], SFL [131], FedPoiRec [30],
+PrefFedPOI [132], PREFER [133], Dong et al. [134], DMF [135], NRDL [136] , modelnode-paradigms, text width=40em] ] ] [ Security (Sec. VI), leaf-others,text width=8em [ System Vulnerability
+Analysis, leaf-others, text width=9.5em [FedRecAttack [137], Ghost riders [138] , LOKI [139], Yuan et al.
+[140], PTIA [141] , modelnode-others, text width=40em] ] [ User Privacy Protection, leaf-others, text width=9.5em [DMF [135], Li et al.
+[142], PriRec [143], Liu et al.
+[144], Huo et al.
+[145], FedPoiRec [30], modelnode-others, text width=40em] ] ] ]
+Figure 1: Taxonomy of existing studies for POI recommendation in terms of models, architecture and security
 III Taxonomy
-This paper distinguishes itself from existing surveys on POI recommendation by adopting a holistic approach that emphasizes three critical and interrelated aspects: models, architectures, and security.
-This tripartite framework, as illustrated in Fig.
-2, serves as a comprehensive taxonomy for systematically analyzing and comparing various POI recommendation systems.
-To further contextualize the research landscape, we present the distribution of existing studies across these three categories in Fig.
-1 (a).
-The figure underscores a clear trend: methods for improving POI recommendation models dominate the field, comprising over 75% of the total research efforts.
-This focus reflects the ongoing challenge of enhancing accuracy, personalization, and context-awareness in POI recommendations.
-On the other hand, studies addressing security issues in POI recommendations remain underrepresented, making up approximately 8% of the existing research.
-This highlights a significant gap in the literature, given the increasing importance of safeguarding user privacy, defending against adversarial attacks, and ensuring data integrity within location-based services.
-In contrast, research on architectures, which bridges the gap between models and real-world deployment, occupies the remaining portion, indicating its crucial yet often overlooked role in ensuring the scalability, efficiency, and responsiveness of POI recommendation.
-In the rest of this section, the introduction of the taxonomy will unfold along these three dimensions.
+This paper distinguishes itself from existing surveys on POI recommendation by adopting a holistic approach that focuses on three critical, interrelated aspects: models, architectures, and security.
+This tripartite framework serves as a comprehensive taxonomy for analyzing and comparing various POI recommendation systems.
+To contextualize the research landscape, Fig.
+1 illustrates the framework, while Fig.
+2 (a) presents the distribution of existing studies across these three categories.
+Research on POI recommendation models dominates the field, accounting for over 75% of studies, reflecting ongoing challenges in improving accuracy, personalization, and context-awareness.
+In contrast, security-related research remains limited, comprising around 8% of the literature, underscoring gaps in addressing user privacy, adversarial threats, and data integrity in location-based services.
+Research on architectures, which ensures scalability and efficiency in real-world deployments, occupies the remaining portion, emphasizing its critical yet often overlooked role.
+To further illustrate the evolution of methodologies in POI recommendation, Fig.
+2 (b) provides a structured timeline, where key techniques are arranged chronologically from bottom to top.
+This visualization highlights the shift from traditional models such as LDA and MF to more advanced techniques, including LSTMs, GNNs, transformers, and large language models (LLMs).
+Additionally, it categorizes architectures into centralized, decentralized, and federated-learning-based approaches, while security-related concerns are mapped to areas such as user privacy protection and system vulnerability analysis.
+This progression underscores the increasing sophistication and diversification of POI recommendation research over time.
 Refer to caption Refer to caption
@@
 III-A Models
-Building on previous POI recommendation surveys [38, 39], which primarily reviewed studies based on traditional non-deep learning and early deep learning methods, there remains a need to address the advances brought by more recent techniques.
-With the emergence of novel approaches such as diffusion models, self-supervised learning, and large language models, a growing body of research has evolved that is not yet covered in existing surveys.
-To fill this gap, we provide an extensive review of models spanning from traditional methods to cutting-edge techniques, as illustrated in Fig.
-2.
-To further clarify the progression of these methodologies, we present a visual depiction of their structural evolution in Fig.
-1 (b), where the labels are arranged chronologically from bottom to top.
-This illustration highlights how the research landscape has shifted over time, with an increasing focus on more sophisticated and advanced models.
+Building on previous POI recommendation surveys [146, 147], which primarily reviewed studies based on traditional non-deep learning and early deep learning methods, we address the advancements brought by more recent techniques.
+With the rise of novel approaches such as diffusion models, self-supervised learning, and large language models, a growing body of research remains uncovered in existing surveys.
+To bridge this gap, we provide an extensive review of models from traditional methods to cutting-edge techniques.
 Detailed explanations of each method, along with the corresponding preliminaries, are provided in Section IV.
@@
 It can also be used to pretrain the recommendation model by encouraging the consistency between augmentations of the same user/item while minimizing similarities between augmentations of different users/items.
-III-A4 Generative Model-Based
+III-A4 Generative
 Generative models use machine learning to discover patterns in data and generate new content.
@@
 VAEs encode data into a latent space and decode it to generate new instances.
-LLMs, renowned for their ability to comprehend and generate human-like text, have emerged as a transformative force in recommendation systems, including POI recommendations [46, 47].
+Large language models, renowned for their ability to comprehend and generate human-like text, have emerged as a transformative force in recommendation systems, including POI recommendations [112, 113].
 Traditionally, recommendation models relied on structured data like user-item interactions, but LLMs introduce a paradigm shift by incorporating rich, unstructured textual data, including user reviews, POI descriptions, and contextual insights such as time, location, and trends.
@@
 $p(r_{i}|u_{j},c_{i},\theta)$ denotes the probability of generating recommendation $r_{i}$ given user profile $u_{j}$, context $c$, and parameters $\theta$.
-DMs, which learn to gradually add noise to data and reverse the process to generate new samples, can be applied to POI recommendation by generating plausible user-POI interaction patterns.
+Diffusion models, which learn to gradually add noise to data and reverse the process to generate new samples, can be applied to POI recommendation by generating plausible user-POI interaction patterns.
 They can create new user-POI interaction data, effectively filling in missing entries in the interaction matrix.
@@
 $D_{KL}(.)$ is Kullback-Leibler divergence, measuring the difference between two probability distributions.
-VAEs [48, 49] are generative models that learn to represent data in a compressed latent space.
+Variational autoencoders [151, 152] are generative models that learn to represent data in a compressed latent space.
 They consist of an encoder mapping input data to a distribution in this latent space and a decoder reconstructing data from samples drawn from that distribution.
@@
 This formulation assumes access to all available data in one place, which, while advantageous for performance optimization, amplifies privacy risks and imposes significant computational overhead.
-III-B2 Decentralized-Based
-Decentralized-based structures shift the processing burden from centralized servers to users’ individual devices (e.g., smartphones) [54].
+III-B2 Decentralized
+Decentralized architectures shift the processing burden from centralized servers to users’ individual devices (e.g., smartphones) [124].
 In this approach, recommendation models are deployed locally on the user’s device, where data processing and model inference occur without the need to continuously communicate with a central server.
@@
 $\mathbf{Enc}(x)\otimes\mathbf{Enc}(y)$ $=\mathbf{Enc}(x\cdot y)$
-where $\mathbf{Enc}(\cdot)$ denotes the encryption algorithm under a specific key, $\oplus$,$\otimes$ , ⊗ represent addition and multiplication over ciphertext, +,+$\cdot$+ , ⋅ represent addition and multiplication over plaintext, and $x,y$ are the input data.
+where $\mathbf{Enc}(\cdot)$ denotes the encryption algorithm under a specific key, $\oplus$ and $\otimes$ represent addition and multiplication over ciphertext, $+$ and $\cdot$ represent addition and multiplication over plaintext, and $x,y$ are the input data.
 This property allows computations to be performed on encrypted data, producing encrypted results that can be decrypted to obtain the final output.
 These methods facilitate data analysis without compromising the confidentiality of personal information, allowing systems to generate accurate recommendations while maintaining user trust and privacy.
-{forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=center, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [POI Recommendation,leaf-head, ver [ Models (Sec.
-IV), leaf-task,text width=7em [ Latent Factor Model-Based, leaf-task10, text width=11.5em [LDA-Based, leaf-task8, text width=7.5em[LCARS [64, 16], JIM [65], Geo-SAGE [66], TRM
- [65], LA-LDA [67], TRM+ [68], SPORE [69], ST-LDA [70],
-UCGT [71], LSARS [72], ST-SAGE [73], TPM [74], modelnode-task8,text width=29em]] [MF-Based, leaf-task9, text width=7.5em[GeoMF [75], GeoMF++ [76], LGLMF [77], SSTPMF [78],
-SQPMF [79], modelnode-task9,text width=29em]] ] [ Classic NN-Based, leaf-taska, text width=9.5em [LSTM-Based, leaf-task1, text width=9.5em [ST-LSTM [80], LSTM-S [81], PS-LSTM [82], STMLA [83], LSMA
- [84], AT-LSTM [85], ATSD-GRU [86], DLAN [87],modelnode-task1, text width=29em] ] [Transformer-Based, leaf-task2, text width=9.5em [TLR-M [41], GETNext [88], CARAN [89], JANICP [90],
-Li et al. [91], STUIC-SAN [92], AMACF [93], CHA [94],
-MVSAN [95], HAT [96], STAR-HiT [97], CAFPR [98],
-FPG [99], TGAT [100], MobGT [101], POIBERT [102],
-AutoMTN [103], CCDSA [104], Liu  et al. [105], TDGCN [106],
-TGAT [100], MGAN [107] , STA-TCN [108] , Xia et al. [109],
-STTF-Recommender [110] , BayMAN [111], Kumar et al. [112],
-ImNext [113] , BSA-ST-Rec [114] , Wu et al.
-[115] , ROTAN [116],modelnode-task2, text width=29em] ] [GNN-Based, leaf-task3, text width=9.5em [STGCN [117], Dynaposgnn [20] , ASGNN [21] , GNN-POI [118] ,
-ADQ-GNN [22] , GN-GCN [119] , ATST-GGNN [120] , Zhao et al.
- [121] , HS-GAT [122] , PROG-HGNN [28] , CGNN-PRRG [123] ,
-HKGNN [124],modelnode-task3, text width=29em] ] ] [ SSL-Based Models, leaf-task6, text width=9.5em [SML [125], S2GRec [126], SSTGL [127], SelfTrip [128], GSBPL [44], LSPSL [129],
-Gao et al.
-[130], Wang et al.
-[131], SLS-REC [132], CLLP [45], , modelnode-task6, text width=40em] ] [ Generative-Based, leaf-taska, text width=9.5em [ LLM-Based, leaf-task5, text width=9.5em [LLMmove [46], Li et al.
-[47], modelnode-task5,text width=29em] ] [ DM-Based, leaf-task7, text width=9.5em [Diff-POI [133], DSDRec [134] , Diff-DGMM [135] , DCPR [136], modelnode-task7,text width=29em] ] [VAE-Based, leaf-task4, text width=9.5em [WaPOIR [137], Li et al.
-[138], modelnode-task4, text width=29em] ] ] ] [ Architectures
-(Sec.
-V), leaf-paradigms,text width=7em [ Centralized, leaf-paradigms, text width=9.5em [ GeoIE [51], SH-CDL [50], GETnext [88], STGN [139], AGRAN [53] , modelnode-paradigms, text width=40em] ] [ Decentralized-Based, leaf-paradigms, text width=9.5em [ MAC [54], DARD [140], LLRec [141], DCLR [27], modelnode-paradigms, text width=40em] ] [ FL-Based, leaf-paradigms, text width=9.5em [ FL&PP-POI [142], Huang et al.
-[143],CPF-POI [144], SCFL [145], SFL [146], FedPoiRec [29],
-PrefFedPOI [147], PREFER [148], Dong et al.
-[149], DMF [60], NRDL [150] , modelnode-paradigms, text width=40em] ] ] [ Security (Sec.
-VI), leaf-others,text width=7em [ Data Integrity Threats, leaf-others, text width=9.5em [FedRecAttack [56], Ghost riders [58] , LOKI [57], Yuan et al.
-[151], PTIA [59] , modelnode-others, text width=40em] ] [ User Privacy Protection, leaf-others, text width=9.5em [DMF [152], Li et al.
-[62], PriRec [61], Liu et al.
-[63], Huo et al.
-[153], FedPoiRec [29], modelnode-others, text width=40em] ] ] ]
-Figure 2: Taxonomy of existing studies for POI recommendation in terms of models, architecture and security
 IV Models
@@
 LSTM networks have become indispensable for POI recommendation [56, 153, 57, 59, 60], effectively modeling the sequential dependencies inherent in user check-in data to predict future preferences.
-Earlier studies [80, 52, 81] adopt LSTM to model POI sequences.
-Zhang et al.
-[80] propose a novel approach to learning user preferences by introducing the Discrete Ranking-based Matrix Factorization (DRMF) algorithm.
-DRMF leverages pairwise user preferences, framing the learning process as a binary quadratic programming problem to derive binary codes.
-To address the challenges of non-convexity and binary constraints inherent in this formulation, they introduce a self-paced learning strategy.
-This strategy gradually incorporates pairwise preferences, starting with simpler relationships and progressing to more complex ones, thereby enhancing the optimization process.
-Nevertheless, Zhao et al.
+Earlier studies [56, 153, 57] adopt LSTMs to model POI sequences.
+Zhao et al.
 [153] address the limitations of traditional RNNs in capturing the crucial influence of time and distance between consecutive check-ins.
 Their proposed STLSTM model integrates spatio-temporal context, short-term spatio-temporal gating, and long-term spatio-temporal gating.
-Another earlier study [81] introduces an approach that incorporates temporal information to understand the semantic connections between user actions.
-This approach leverages a semantic correlational graph to calculate semantic sequential correlation, factoring in time intervals.
-It then integrates this semantic information into a modified LSTM framework, incorporating two additional semantic gates.
+Another study by Zhan et al.
+[57] incorporates temporal information to understand the semantic relationships between user actions.
+They leverage a semantic correlational graph to calculate semantic sequential correlation, factoring in time intervals, and integrate this information into a modified LSTM framework with two additional semantic gates.
 This enhanced model captures users’ sequential behaviors and long- and short-term interests within a semantic context.
-Finally, this approach proposes user clustering to improve recommendation accuracy by grouping users based on their semantic preferences.
+It also introduces user clustering to improve recommendation accuracy by grouping users based on their semantic preferences.
 Recent studies [60, 63] based on LSTM mainly aim to capture dynamic dependencies in long-term POI sequences.
 Yang et al.
-[84] tackle the challenge of effectively capturing the dynamic interplay between a user’s long-term preferences and their more immediate, context-dependent interests, namely AMACF.
+[60] tackle the challenge of effectively capturing the dynamic interplay between a user’s long-term preferences and their more immediate, context-dependent interests with AMACF.
 It achieves this by dynamically weighting input sequences and seamlessly fusing long- and short-term signals.
@@
 This parallel approach addresses the limitations of RNN-based methods, which struggle to establish long-term dependencies between sequences.
-Furthermore, DLAN integrates a user similarity weighting layer to quantify the influence of social relationships between users on the target user’s preferences.
-This layer leverages social connections to improve the accuracy of recommendations by incorporating insights from similar users.
-Beyond these examples, other studies such as Zhong et al.
+Furthermore, DLAN integrates a user similarity weighting layer to quantify the influence of social relationships between users on the target user’s preferences, leveraging social connections to improve the accuracy of recommendations.
+Beyond these studies, others such as Zhong et al.
 [58] and Zhang et al.
-[83] also explore innovative ways to leverage LSTMs for POI recommendation, further demonstrating the versatility and effectiveness of this approach.
-These studies highlight the ongoing evolution of LSTM-based methods for POI recommendation.
-Researchers continue to develop sophisticated techniques to capture the complexities of user behavior, ultimately leading to more accurate, personalized, and context-aware recommendations.
+[59] explore innovative ways to leverage LSTMs for POI recommendation, further demonstrating the versatility and effectiveness of this approach.
+These studies highlight the ongoing evolution of LSTM-based methods for POI recommendation, as researchers continue to develop sophisticated techniques to capture the complexities of user behavior, ultimately leading to more accurate, personalized, and context-aware recommendations.
 IV-B2 Transformer
@@
 Each node $v\in V$ is associated with a feature vector $h^{\prime}_{v}$.
-GNNs utilize the structure of the graph, along with these node features $h^{\prime}_{v}$, to learn a representation vector for each node, $h_{v}$, or for the entire graph, $h_{G}$.
+GNNs utilize the structure of the graph along with these node features $h^{\prime}_{v}$ to learn a representation vector for each node, $h_{v}$, or for the entire graph, $h_{G}$.
 Modern GNN architectures typically employ a neighborhood aggregation strategy, where a node’s representation is iteratively refined by aggregating the representations of its neighbors [161].
@@
 where $h_{v}^{(k)}$ is the feature vector of node $v$ at the $k$-th iteration, with $h_{v}^{(0)}$ initialized to $X_{v}$.
-The set $\mathcal{N}(v)$ denotes the nodes adjacent to $v$.
-The functions $\mathsf{Aggregate}^{(k)}(\cdot)$ and $\mathsf{Combine}^{(k)}(\cdot)$ are crucial to the design of GNNs.
+The set $\mathcal{N}(v)$ denotes the neighbors of node $v$.
+The functions $\mathsf{Aggregate}^{(k)}(\cdot)$ and $\mathsf{Combine}^{(k)}(\cdot)$ are crucial components of GNN design.
 GNNs have emerged as a powerful tool for POI recommendation [94, 21, 22, 95, 97, 100, 101], effectively capturing complex relationships within location-based data.
-Earlier studies [117, 20, 21, 118, 120] mainly emphasize capturing spatial and temporal dynamics.
+Earlier studies [94, 21, 22, 95, 97] focus on capturing spatial and temporal dynamics.
 Specifically, Han et al.
-[117] focuses on capturing both spatial and temporal dynamics within user check-in data.
-First, it constructs a multigraph to represent various contextual factors and their relationships.
-Then, it employs a time-aware sampling strategy to capture the evolving influence of neighboring nodes.
-Finally, it learns node representations that change over time, reflecting the dynamic nature of user preferences and POI popularity.
-Another earlier study [20] proposes a novel next POI recommendation model called DynaPosGNN, which incorporates specific arrival times into offline user activities.
-Unlike traditional methods that solely rely on user check-in history, DynaPosGNN analyzes the correlation between arrival time and two spatial dynamic graphs: the “User-POI graph” and the “POI-POI graph”.
-By integrating temporal information, DynaPosGNN provides a more nuanced understanding of user behavior, considering both the locations visited and the timing of those visits, leading to potentially more accurate recommendations.
-In comparison, ASGNN [21] incorporates a personalized hierarchical attention network to exploit complex correlations between users and POIs in check-in sequences.
-This allows the model to capture both long-term and short-term user preferences, providing a more discriminative understanding of user behavior.
-Finally, ASGNN leverages these learned user preferences to predict the next POI, taking into account both long-term and short-term behavioral patterns.
-Zhang et al.
-[118] introduce GNN-POI, a POI recommendation framework that leverages the power of GNNs to learn node representations from both node information and topological structure, thus enhancing POI recommendation accuracy.
-GNNs have demonstrated exceptional capacity for modeling complex relationships within graph structures, making them well-suited for capturing the intricate connections between users and POIs in LBSNs.
+[94] captures both spatial and temporal dynamics in user check-in data.
+It constructs a multigraph to represent contextual factors and their relationships, employs a time-aware sampling strategy to account for the evolving influence of neighboring nodes, and learns time-dependent node representations to reflect changing user preferences and POI popularity.
+Another study [21] introduces DynaPosGNN, a next POI recommendation model that integrates arrival times into user activities.
+Unlike traditional methods that rely solely on check-in history, DynaPosGNN analyzes the correlation between arrival time and two spatial dynamic graphs—the ”User-POI graph” and the ”POI-POI graph”—providing a more nuanced understanding of user behavior.
+ASGNN [22] incorporates a personalized hierarchical attention network to capture complex correlations between users and POIs in check-in sequences.
+This enables the model to capture both long-term and short-term user preferences, enhancing its ability to predict the next POI.
+Finally, Zhang et al.
+[95] introduce GNN-POI, a POI recommendation framework that utilizes GNNs to learn node representations from both node information and topological structure, improving POI recommendation accuracy.
+GNNs’ exceptional capacity to model complex relationships makes them well-suited for capturing the intricate connections between users and POIs in LBSNs.
 Refer to caption Refer to caption
@@
 Figure 4: GNN and DM model structures.
-Recent studies [124, 123] based on GNN focus on adopting GNN to capture side information.
+Recent studies [101, 100] based on GNN focus on using GNNs to capture side information.
 Liu et al.
-[123] propose an approach to learning similar users’ POI transfer preferences using a session-aware GNN.
-This method aims to understand how users’ preferences for POIs evolve over time and how those preferences can be transferred to similar users.
-By contrast, HKGNN [124] leverages the power of hypergraphs and side information to enhance POI recommendations.
-The key details include utilizing a hypergraph to represent the complex relationships between users, POIs, and other relevant entities, employing a hypergraph neural network to learn from the rich structural information encoded in the HKG, and incorporating side information, such as POI categories and user demographics, to mitigate data sparsity and enhance recommendation accuracy.
+[100] propose a session-aware GNN to learn POI transfer preferences from similar users.
+This approach aims to understand how users’ POI preferences evolve over time and how these preferences can be transferred to similar users.
+In contrast, HKGNN [101] utilizes hypergraphs and side information to enhance POI recommendations.
+It employs a hypergraph to represent complex relationships between users, POIs, and other entities, a hypergraph neural network to learn from the structural information encoded in the hypergraph, and side information such as POI categories and user demographics to address data sparsity and improve recommendation accuracy.
 IV-C Self-Supervised Learning-Based
 Self-supervised learning (SSL) [149] is a versatile framework that leverages unsupervised data by creating surrogate tasks, known as pretext tasks, to help models learn meaningful representations.
-These tasks are specifically designed to guide the learning process towards capturing useful features from the data.
+These tasks are designed to guide the learning process towards capturing useful features from the data.
 SSL methods are generally categorized into two approaches: generative and contrastive [162].
 Generative SSL aims to create or reconstruct data, focusing on learning detailed representations, while contrastive SSL emphasizes comparing data samples by bringing similar pairs closer in the latent space and pushing dissimilar pairs apart.
-The SSL process typically begins with defining a pretext task, which does not require labeled data but helps in learning valuable representations.
-In contrastive learning, the pretext task might involve differentiating between augmented versions of the same image (positive pairs) and different images (negative pairs).
+The SSL process typically begins by defining a pretext task, which does not require labeled data but helps in learning valuable representations.
+In contrastive learning, the pretext task involves differentiating between augmented versions of the same image (positive pairs) and different images (negative pairs).
 The model is trained by minimizing the following loss function [163]:
 $\mathcal{L}_{\text{SSL}}=-\log\frac{\mathsf{Exp}(\mathsf{Sim}(\mathsf{pair}_{i},\mathsf{pair}_{j})/\tau)}{\sum_{k=1}^{2N}\mathbf{1}_{[k\neq i]}\mathsf{Exp}(\mathsf{Sim}(\mathsf{pair}_{i},\mathsf{pair}_{k})/\tau)}$ (10)
-where $\mathbf{1}_{[k\neq i]}\in\{0,1\}$ is an indicator function that equals 1 only if $k\neq i$, $\mathsf{pair}_{i}$ and $\mathsf{pair}_{j}$ are the feature representations of a positive pair, $\mathsf{Sim}(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is a temperature scaling parameter, and $N$ represents the number of data points.
-Self-supervised learning has emerged as a promising paradigm in POI recommendation [163].
+where $\mathbf{1}_{[k\neq i]}\in\{0,1\}$ is an indicator function that equals 1 only if $k\neq i$, $\mathsf{pair}_{i}$ and $\mathsf{pair}_{j}$ are the feature representations of a positive pair, $\mathsf{Sim}(\cdot,\cdot)$ denotes cosine similarity, $\tau$ is a temperature scaling parameter, and $N$ represents the number of data points.
+SSL has emerged as a promising paradigm in POI recommendation [164].
 By leveraging patterns and structures within data, self-supervised techniques enable models to learn meaningful representations without explicit labels.
@@
 Self-supervised POI recommendation systems [102, 103, 105, 107, 106, 109, 110] can be divided into two categories.
-The first category [125, 126, 128] focuses on creating augmented views through human design.
+The first focuses on creating augmented views through human design [102, 103, 105].
 Zhou et al.
-[125] introduce a Self-supervised Mobility Learning (SML) framework to encode human mobility semantics and enhance location-based tasks.
+[102] introduced a Self-supervised Mobility Learning (SML) framework to enhance location-based tasks by encoding human mobility semantics.
 SML addresses challenges posed by sparse and noisy mobility data by leveraging spatio-temporal contexts and augmented traces.
-SML generates contrastive views via traditional ways like sampling trajectory points.
-Another work [126] introduces a paradigm that contrastively learns trajectory representations from geolocation and temporal transitions via random sampling POI points to obtain contrastive views, enhancing the model’s understanding of mobility patterns.
-At its core is a Graph-enhanced Self-attentive layer, which uncovers transitional dependencies among POIs and captures temporal interests.
-Lastly, SelfTrip [128] focuses on trip recommendation using a two-step contrastive learning mechanism to learn nuanced representations of POIs and itineraries based on generated views via random walk, further enhanced by four innovative trip augmentation methods.
-Nonetheless, recent studies [129, 44, 131, 132] on the self-supervised paradigm focus on creating augmented views automatically.
-While many existing studies prioritize final POI recommendation performance, they often overlook the crucial connection between POI sequences and contextual information.
-This can lead to limited data representations and less accurate recommendations.
+It generates contrastive views through methods such as sampling trajectory points.
+Another work [103] introduces a paradigm that learns trajectory representations via random sampling of POI points, enhancing the model’s understanding of mobility patterns.
+A Graph-enhanced Self-attentive layer uncovers transitional dependencies among POIs and captures temporal interests.
+Lastly, SelfTrip [105] focuses on trip recommendation using a two-step contrastive learning mechanism to learn nuanced representations of POIs and itineraries based on augmented views from random walks, further enhanced by four innovative trip augmentation methods.
+Recent studies [107, 106, 109, 110] focus on automatically creating augmented views.
+While many studies prioritize POI recommendation performance, they often overlook the connection between POI sequences and contextual information, leading to limited data representations and less accurate recommendations.
 To address this, Jiang et al.
-[129] propose LSPSL, a unified attention framework for next POI recommendation that leverages self-supervised learning to model users’ long- and short-term preferences.
-LSPSL utilizes a self-attention network and two self-supervised optimization objectives to deeply explore the inherent relationships between POI sequences and contextual information and generate contrastive views automately during a pre-training phase.
-This process strengthens the data representations, allowing the model to better understand the nuances of user behavior.
-GSBPL [44] has been introduced for the domain of next-POI recommendation.
-GSBPL utilizes a pair of graph data augmentation techniques based on modeling user behaviour patterns via GNNs, to create augmented trajectory graphs, enabling the effective modeling of implicit behavioral patterns.
-To acquire high-order representations of trajectory graphs, GSBPL incorporates a graph preference representation encoder that utilizes geographical and social context.
-By applying contrastive learning, GSBPL captures implicit behavior patterns within these representations.
-To address the limitations of existing approaches that neglect the complex higher-order relationships between user-POI dynamics over time and fail to account for the varying importance of POIs, leading to an inability to effectively capture geographical influence imbalances, Fu et al.
-[132] introduce SLS-REC, a novel self-supervised long- and short-term model for POI recommendation.
-SLS-REC leverages a spatio-temporal attention hypergraph neural network to capture the spatial dependencies and temporal evolution of the short-term dynamic interests of users and creates contrastive views automately.
-IV-D Generative Model-Based
+[107] propose LSPSL, a unified attention framework for next-POI recommendation that leverages self-supervised learning to model long- and short-term preferences.
+LSPSL uses a self-attention network and two self-supervised optimization objectives to explore relationships between POI sequences and contextual information, generating contrastive views automatically during pre-training.
+GSBPL [106] introduces graph-based data augmentation techniques for next-POI recommendation, modeling user behavior patterns via GNNs to create augmented trajectory graphs.
+By applying contrastive learning, GSBPL captures implicit behavioral patterns.
+To address limitations in capturing higher-order relationships and the varying importance of POIs, Fu et al.
+[110] introduce SLS-REC, a self-supervised model for POI recommendation.
+SLS-REC leverages a spatio-temporal attention hypergraph neural network to capture spatial dependencies and the temporal evolution of short-term dynamic user interests while automatically creating contrastive views.
+IV-D Generative
 IV-D1 Large Language Model-Based
@@
 The decoder reconstructs check-in patterns, and a classifier associates representations with users, facilitating a structured approach to user modeling.
-TABLE II: Model comparison in terms of effectiveness, efficiency, scalability and explainability
+IV-E Discussion of Models
+In this section, we compare the model types discussed earlier: Latent Factor Models (LFMs), Classic Neural Networks (NNs), Self-Supervised Learning (SSL), and Generative Models (Gen-Models), focusing on model effectiveness, inference efficiency, training scalability, and result explainability, as summarized in Table III.
+TABLE III: Comparison of models based on effectiveness, efficiency, scalability, and explainability
 Criterion LFMs NNs SSL Gen-Models
-Model Effectiveness Moderate High High High
-Inference Efficiency High Varies Varies Low
-Training Scalability High Moderate Moderate Low
-Result Explainability Moderate Low Moderate High
-IV-E Discussion of Models
-In this part, we further compare the discussed model types: Latent Factor Models (LFMs), Classic Neural Networks (NNs), Self-Supervised Learning (SSL), and Generative Models (Gen-Models) from the aspects of model effectiveness, inference efficiency, training scalability, result explainability, shown in Table II.
-The comparison analysis is shown as:
-Model effectiveness:
-  • LFMs: LFMS moderately captures user-POI interactions using latent features for comparable recommendations.
-  • NNs: NNs potentially achieve higher effectiveness by learning complex non-linear relationships between users and POIs, leading to more nuanced recommendations.
-  • SSL: The effectiveness of SSL can be high, particularly when labeled data is scarce, by learning robust representations from unlabeled data.
-  • Gen-Models: Gen-Models obtain high potential effectiveness due to their ability to generate diverse POI recommendations based on learned data distributions.
-Inference efficiency:
-  • LFMs: LFMS have high prediction efficiency, especially for smaller datasets, resulting in small prediction time.
-  • NNs: Efficiency of NNs varies greatly depending on the network architecture and size; larger, more complex networks are less efficient.
-  • SSL: Efficiency depends on the specific method employed; some approaches are more efficient than others.
-  • Gen-Models: The inference efficiency of Gen-Models lower than other kinds of models due to their high model complexity and large model volume.
-Training scalability:
-  • LFMs: Scalability of training is higher than other models due to its simple model structure.
-  • NNs: Training scalability is moderate; it’s limited only when the computational resources required for training complex networks become insufficient.
-  • SSL: SSL often exhibits similar training scalability as NNs, especially methods using simpler architectures.
-  • Gen-Models: Scalability is a major challenge; training scalability becomes computationally expensive with increasing data size.
-Result explainability:
-  • LFMs: Results of LFMs offer moderate explainability; latent factors provide some insights into user preferences.
-  • NNs: Explainability is generally low; understanding results of complex neural networks is difficult.
-  • SSL: Explainability is moderate; interpreting the learned representations requires further analysis.
-  • Gen-Models: Explainability is higher than results of other models; understanding generated recommendation results often necessitates analyzing the latent space.
+Effectiveness Moderate High High High
+Efficiency High Varies Varies Low
+Scalability High Moderate Moderate Low
+Explainability Moderate Low Moderate High
+Firstly, in terms of model effectiveness, LFMs provide moderate performance by using latent features to generate reasonable recommendations, though with limited depth.
+NNs, on the other hand, excel at capturing complex relationships, leading to more nuanced recommendations.
+SSL is particularly effective when labeled data is scarce, as it learns robust representations from unlabeled data.
+Gen-Models offer high effectiveness, especially when diverse recommendations are needed, due to their generative nature.
+Secondly, regarding inference efficiency, LFMs are highly efficient, particularly with smaller datasets.
+NNs’ efficiency depends on the network architecture, with larger models resulting in slower inference.
+The efficiency of SSL varies depending on the method used, with some being computationally expensive.
+Gen-Models tend to be less efficient due to their complexity and large parameter volumes, which result in longer inference times.
+Additionally, in terms of training scalability, LFMs are highly scalable due to their simple structure, enabling efficient handling of large datasets.
+NNs and SSL have moderate scalability, as their training becomes computationally intensive with larger datasets.
+Gen-Models face significant scalability challenges, as training them becomes prohibitively expensive with the increasing data size.
+Finally, in terms of result explainability, LFMs offer moderate transparency, providing some insight into user preferences through latent factors.
+NNs generally lack explainability due to their complex architectures, making it difficult to interpret their decisions.
+SSL provides moderate explainability, though interpreting the learned representations often requires further analysis.
+Gen-Models provide higher explainability than the other models, but understanding them fully necessitates exploring the latent space, which can be complex.
 V Architecture
@@
 V-A Centralized Architecture
-Centralized architecture serves as a robust, scalable, and cost-effective foundation for building and deploying recommendation systems.
-In this model, centralized data centers and servers provide a consolidated platform for managing large-scale computational resources.
-By centralizing these resources, enterprises can swiftly respond to changes in demand, reduce overhead, and streamline system maintenance.
-A key benefit of this architecture is its ability to centralize data processing and storage, allowing for more efficient management of data flow and integration across systems.
-The flexibility of centralized architecture can be expressed through the following equation:
+Centralized architecture provides a scalable, cost-effective foundation for recommendation systems by using centralized data centers to manage computational resources.
+This model allows enterprises to respond quickly to demand changes, reduce overhead, and streamline maintenance.
+Centralizing data processing and storage enhances efficiency in managing data flow and system integration.
+The flexibility of centralized architecture is expressed by the following equation:
 […]
 Scalability+Cost Efficiency+AccessibilityInitial Hardware Investments+Physical $\frac{\text{Scalability}+\text{Cost Efficiency}+\text{Accessibility}}{\text{Initial Hardware Investments}+\text{Physical Constraints}}$
-This equation highlights how centralized systems optimize scalability, cost-effectiveness, and resource availability, enabling rapid adaptation without requiring extensive upfront hardware investments.
+This equation illustrates how centralized systems optimize scalability, cost-effectiveness, and resource availability with minimal upfront hardware investments.
 We explore studies [120, 121, 65, 122, 123] on POI recommendation systems developed using centralized architecture.
 For instance, Wang et al.
-[51] leverage centralized computing to model geographical influences in POI recommendations with GeoIE.
+[120] leveraged centralized computing to model geographical influences in POI recommendations with GeoIE.
 Centralized servers enable real-time processing of geographical data, enhancing the efficiency of recommendations by factoring in geo-influence, geo-susceptibility, and physical distance.
@@
 However, as the complexity of these systems increases, continuous innovation in defense mechanisms is essential to safeguard data integrity and user privacy within centralized frameworks.
-V-B Decentralized-based Architecture
+V-B Decentralized Architecture
 The architecture of decentralized methodologies prioritizes the maximal utilization of local storage and processing capabilities, thereby conserving bandwidth and curtailing cloud storage expenses.
@@
 where the decentralized efficiency metric integrates local processing power, storage capacity, bandwidth consumption, and cloud storage expenses to gauge the overall effectiveness of decentralized computing solutions in comparison to cloud-based alternatives.
-TABLE III: Related works in different application sub-tasks.
-We provide detailed information on the names of journals/conferences, the specific generative techniques used, the particular application sub-tasks addressed, and the datasets used for evaluation.
+In this part, we illustrate the decentralized architecture for POI recommendation.
+While centralized architectures excel in accuracy, they require substantial computational resources during training.
+To address these challenges, researchers have introduced decentralized architectures like [124, 28, 125].
+Specifically, Long et al.
+[124] propose Model-Agnostic Collaborative Learning (MAC), which allows users to tailor model architectures, such as adjusting dimensions and the number of hidden layers.
+To mitigate the scarcity of user data on devices, MAC [124] pre-determines collaborators based on physical proximity, preferences, and social connections.
+It integrates insights from collaborators using a knowledge distillation approach, maximizing mutual information, and selectively sh-
+TABLE IV: Related works in various application sub-tasks. We provide details on journal/conference names, generative techniques used, specific sub-tasks addressed, and evaluation datasets.
 Models Source Technique Sub-task Datasets
@@
 LDA-based Models
-LCLCARS [64] KDD 2013 LDA Location-Content-Aware Recommender Foursquare, and Douban Event
-LCARS [16] TOIS 2014 LDA Spatial Item Recommender Douban Event, and Foursquare
+LCLCARS [40] KDD 2013 LDA Location-Content-Aware Recommendation Foursquare, and Douban Event
+LCARS [17] TOIS 2014 LDA Spatial Item Recommendation Douban Event, and Foursquare
 JIM [41] CIKM 2015 LDA POI Recommendation Foursquare, and Twitter
 Geo-SAGE [42] KDD 2015 LDA Spatial Item Recommendation Foursquare, and Twitter
-TRM [65] ACM MM 2015 LDA POI Recommendation Foursquare, and Twitter
+TRM [41] CIKM 2015 LDA POI Recommendation Foursquare, and Twitter
 LA-LDA [43] TKDD 2015 LDA Location-based Recommendation MovieLens, Gowalla, and Douban Event
 TRM+ [44] TOIS 2016 LDA POI Recommendation Foursquare, and Twitter
-SPORE [69] ICDE 2016 LDA Spatial Item Recommendation Foursquare, Twitter [168], and Synthetic Dataset
+SPORE [45] ICDE 2016 LDA Spatial Item Recommendation Foursquare, Twitter, and Synthetic Dataset
 ST-LDA [46] TKDE 2016 LDA POI Recommendation Foursquare, and Gowalla
@@
 ST-SAGE [49] TIST 2017 LDA Spatial Item Recommendation Foursquare, and Twitter
-TPM [74] TIST 2018 LDA Spatial Item Recommendation Foursquare, Twitter [168], and Synthetic Dataset
+TPM [50] TIST 2018 LDA Spatial Item Recommendation Foursquare, Twitter, and Synthetic Dataset
 MF-based Models
@@
 STMLA [59] ICWS 2022 LSTM POI Recommendation Gowalla, Foursquare
-LSMA [84] IJGI 2022 LSTM Next POI Recommendation Foursquare-Charlotte (CHA) [169], Foursquare-NYC
+LSMA [60] IJGI 2022 LSTM Next POI Recommendation Foursquare-Charlotte (CHA), Foursquare-NYC
 AT-LSTM [61] IJSWIS 2023 LSTM POI Recommendation Gowalla, Foursquare
@@
 CHA [71] TOIS 2022 Transformer Next POI Recommendation Foursquare‐NYC, Foursquare-TKY
-HAT [96] IEEE MM 2023 Transformer POI Recommendation BJ, SH, NJ, CD
+HAT [73] TMM 2023 Transformer POI Recommendation BJ, SH, NJ, CD
 STAR-HiT [74] ACM TIS 2023 Transformer POI Recommendation Foursquare‐NYC, Foursquare‐US, Gowalla
-CAFPR [98] Applied SC 2023 Transformer POI Recommendation Foursquare-TKY, CA, Budapest [170], Melbourne [170], Magic k [170]
+CAFPR [75] Applied SC 2023 Transformer POI Recommendation Foursquare-TKY, CA, Budapest, Melbourne, Magic k
 FPG [76] Applied SC 2023 Transformer POI Recommendation Foursquare, Gowalla
@@
 CLLP [111] SIGIR 2024 SSL Next POI Recommendation Foursquare, and Gowalla
+SCL [169] Expert Systems 2025 SSL POI Recommendation Florence, Rome, Pisa, Edinburgh, and Toront
 Generative-based Methods
 LLM-based Methods
-LLMmove [46] Arxiv 2024 LLM POI Recommendation NYC [171], TKY [171]
-Li et al. [47] SIGIR 2024 LLM POI Recommendation NYC [171], TKY [171], Gowalla-CA
+LLMmove [112] CAI 2024 LLM POI Recommendation NYC, TKY
+Li et al. [113] SIGIR 2024 LLM POI Recommendation NYC, TKY, Gowalla-CA
 Diffusion Models (DMs)-based Methods
@@
 Diff-DGMN [116] IOT 2024 Diffusion POI Recommendation IST, JK, SP, NYC, LA
-DCPR [136] arXiv 2024 Diffusion POI Recommendation Foursquare, and Weeplace
+DCPR [117] KDD 2024 Diffusion POI Recommendation Foursquare, and Weeplace
 VAE-based Methods
@@
 Li et al. [119] ICAI 2022 VAE POI Recommendation Gowalla, Foursquare
-In this part, we illustrate the decentralized architecture for POI recommendation.
-While centralized architectures have excelled in accuracy, they require substantial computational resources during the training phase.
-To address these challenges, researchers have introduced decentralized architectures like [54, 27, 140].
-To be specific, Long et al.
-[54] propose an innovative method for decentralized POI recommendations called Model-Agnostic Collaborative Learning for decentralized POI recommendation, which empowers users to tailor their own model architectures, such as adjusting dimensions and the quantity of hidden layers.
-To address the scarcity of user data on devices, MAC [54] introduces a strategy to pre-determine collaborators based on physical proximity, preferences at the category level, and social connections.
-For efficiently and securely integrating insights from these chosen collaborators, MAC implements a knowledge distillation approach with a focus on maximizing mutual information.
-In MAC, rather than exchanging sensitive models or gradients, clients selectively share their informed decisions concerning a predetermined reference dataset.
-However, current collaborative learning systems commonly rely on a single reference dataset shared among all users, which may not cater to individual preferences, potentially hindering knowledge exchange and overall performance.
-Users may struggle to provide relevant insights on items outside their scope of interest, further complicating the collaborative process.
-To bridge this gap, Zheng et al.
-[140] propose the Adaptive Reference Data for Decentralized Collaborative Learning framework.
-DARD tailors adaptive reference datasets for enhanced user collaboration by creating a desensitized public reference pool using transformation and probability data generation techniques.
-For each user, DARD dynamically selects personalized reference data through loss tracking and influence function analysis.
-Individual models are trained using private data and collaboratively with spatial and contextual neighbors.
-During collaboration, users exchange soft decisions based on a blend of their personalized adaptive reference data, fostering more effective collaboration and personalized recommendations.
+aring informed decisions on a reference dataset rather than sensitive models or gradients.
+However, current collaborative learning systems rely on a single shared reference dataset, which may not align with individual preferences and hinder knowledge exchange.
+To address this, Zheng et al.
+[125] propose the Adaptive Reference Data for Decentralized Collaborative Learning (DARD), which creates adaptive reference datasets for improved collaboration.
+DARD uses transformation and probability data generation techniques to create a public reference pool, from which personalized reference data is selected dynamically for each user based on loss tracking and influence function analysis.
+Users train models with private data and collaborate with spatial and contextual neighbors, exchanging soft decisions based on their personalized reference data.
+While both MAC and DARD have merits, they suit different scenarios.
+MAC is effective when a reference dataset is available, such as in cases of public check-in sequences shared under authorization.
+DARD is preferred when no reference data exists, as it creates adaptive datasets for individual preferences.
+Each method excels under specific conditions, and neither method universally outperforms the other.
 V-C Federated Learning-based Architecture
@@
 This formula encapsulates how Federated Learning optimizes data privacy protection, computational resources, bandwidth usage, and model update aggregation to bolster the efficiency of machine learning model training across distributed devices.
-Figure 5: Data poisoning (Left) and model poisoning (Right)
+Figure 5: The centralized, decentralized, and FL-based approaches.
 FL operates on the principle of training machine learning models without requiring access to individual users’ raw private data [127, 128, 129, 130, 131, 30, 132, 133, 134, 136].
-An earlier pioneering work [142] introduces a cross-domain POI framework that incorporates federated learning techniques and privacy safeguards.
-By leveraging data from a related domain in user interest profiling, it aims to mitigate the cold-start issue.
-Additionally, the framework employs federated learning to locally analyze users’ historical data and encrypts latent feature distributions to ensure user privacy while transferring knowledge.
-The method aims to solve challenges like the sensitive and sparse nature of POI data and the cold-start dilemma stemming from data scarcity, which hampers recommendation effectiveness.
-Another earlier work [148] introduces PREFER, a novel federated learning framework optimized for POI recommendation with edge acceleration.
-PREFER divides the recommendation process into distinct segments.
-Initially, for privacy preservation, users develop individualized recommendation models locally and exchange multi-dimensional user-agnostic parameters instead of sharing check-in data.
-Subsequently, to enhance recommendation speed, these decentralized parameters are consolidated on edge servers situated near users (like base stations), as opposed to distant cloud servers.
-Besides, Huang et al.
-[143] present a novel geographical Point of Interest recommendation approach based on federated learning.
-The POI recommendation task is framed as a matrix factorization optimization challenge, with the application of singular value decomposition for matrix decomposition.
-By demonstrating the non-convex nature of the optimization problem, the framework integrates stochastic gradient descent within SVD and devises a FL framework for concurrently addressing the POI recommendation task.
-Within our FL strategy, solely the computed gradient details are transmitted from users to the FL server, while users independently handle their rating and geographic preference data on their personal devices to ensure privacy during communication.
-FedPOIRec [29] is a privacy-focused framework that ensures local data remains on the user’s device while updates are aggregated blindly by a parameter server.
-To enhance personalization, FedPOIRec enables users to exchange learned parameters with their friends, facilitating knowledge transfer.
-To further protect privacy, we introduce a protocol that integrates the preferences of a user’s friends after federated computation.
-This protocol leverages the properties of the CKKS fully homomorphic encryption scheme to ensure privacy during this integration process.
-Figure 6: The centralized, decentralized, and FL-based approaches.
+An earlier work [127] introduces a cross-domain POI framework that incorporates federated learning and privacy safeguards.
+It leverages data from related domains to address the cold-start issue and employs federated learning to analyze users’ historical data locally, encrypting feature distributions to ensure privacy while transferring knowledge.
+This method addresses challenges like sensitive POI data and cold-start problems caused by data scarcity.
+Another work presents PREFER [133], a federated learning framework for POI recommendation with edge acceleration.
+Users develop individualized models locally and exchange multi-dimensional, user-agnostic parameters for privacy.
+These parameters are then consolidated on edge servers, enhancing recommendation speed.
+Huang et al.
+[128] propose a federated learning-based geographical POI recommendation approach, framing the task as a matrix factorization problem solved through singular value decomposition and stochastic gradient descent, with only gradients transmitted for privacy protection.
+In addition, a privacy-focused framework, named FedPOIRec [30], was proposed to integrate the preferences of a user’s friends after the federated computation.
+In this framework, data remains on the user’s device, and updates are aggregated by a parameter server, and it allows users to exchange learned parameters with friends to improve personalization by using a CKKS fully homomorphic encryption scheme to ensure privacy during this process.
+Each method has its strengths depending on the scenario.
+The cross-domain POI framework [127] is useful when related domain data is available to address cold-start issues.
+PREFER [133] excels in environments requiring privacy and fast, localized recommendations using edge servers.
+Huang et al.’s [128] method is ideal when geographical data is crucial, while FedPOIRec [30] is best for applications relying on social connections for personalized recommendations.
+Each approach offers advantages under specific conditions.
 VI Security
 In this section, we provide illustrations on security, including data integrity threats and user privacy protection in POI recommender systems.
-VI-A Data Integrity Threats
+VI-A System Vulnerability Analysis
 Data integrity threats refer to malicious activities that compromise the accuracy, consistency, and trustworthiness of data within systems, particularly those that rely heavily on data-driven technologies.
@@
 The influence of poisoning attacks can be formally represented by manipulating training data, altering the learning process of machine learning models, as shown in the equation $\mathbf{X}_{\mathsf{poisoned}}=\mathbf{X}_{\mathsf{clean}}+\mathbf{A}_{\mathsf{perturb}}$, where $\mathbf{X}_{\mathsf{poisoned}}$ denotes the poisoned data, $\mathbf{X}_{\mathsf{clean}}$ represents the clean data, and $\mathbf{A}_{\mathsf{perturb}}$ signifies the adversarial perturbation introduced during the attack.
+Figure 6: Data poisoning (Left) and model poisoning (Right)
 Rong et al.
@@
 Chen et al.
-[60] propose a decentralized matrix factorization framework for POI recommendation, addressing privacy and computational issues by enabling decentralized training on user devices.
+[135] proposed a decentralized matrix factorization framework for POI recommendation, addressing privacy and computational issues by enabling decentralized training on user devices.
 This approach preserves users’ rating data and enhances recommendation accuracy.
 Furthermore, Chen et al.
-[61] introduce the PriRec framework, which maintains users’ private data on their devices while employing local differential privacy techniques and secure decentralized gradient descent to protect model privacy, achieving comparable recommendation accuracy.
-Figure 7: The workflow of homomorphic encryption
+[143] introduced the PriRec framework, which maintains users’ private data on their devices while employing local differential privacy techniques and secure decentralized gradient descent to protect model privacy, achieving comparable recommendation accuracy.
 Additionally, homomorphic encryption [144, 30] is also commonly used for privacy-preserving POI recommendation.
@@
 Li et al.
-[63] develop a privacy-preserving framework for POI recommendation by utilizing homomorphic encryption.
+[144] developed a privacy-preserving framework for POI recommendation by utilizing homomorphic encryption.
 They address the cold start problem prevalent in POI recommendation systems, which arises due to the sparsity of user-location check-in data.
@@
 By employing offline encryption and parallel computing, Li et al.
-[63] ensure that their protocols efficiently protect the private data of all parties involved in the recommendation process.
+[144] ensured that their protocols efficiently protect the private data of all parties involved in the recommendation process.
 They also prove that their proposed protocols are secure against semi-honest adversaries, and their experiments on both synthetic and real datasets demonstrate the feasibility of privacy-preserving recommendations with acceptable computation and communication costs.
+Figure 7: The workflow of homomorphic encryption
 VII Future Directions
 As POI recommendation systems continue to evolve, vast opportunities for innovation will keep emerging across the three pivotal dimensions: models, architectures, and security.
-By integrating cutting-edge model structures, adopting flexible architectures, and reinforcing security measures, the future of POI recommendation systems are expected to be more adaptive, personalized, and resilient to emerging challenges.
+By integrating cutting-edge model structures, adopting flexible architectures, and reinforcing security measures, the future of POI recommendation systems is expected to be more adaptive, personalized, and resilient to emerging challenges.
 Below, we outline the key future directions that will shape the development of these systems across these essential aspects.
+VII-A Models
 Future POI recommendation models will move beyond traditional and deep learning techniques, incorporating various data modalities and more advanced methodologies to further improve accuracy and personalization.
@@
 By incorporating contextual factors such as location, time, and user-specific constraints, LLM-powered POI recommendation systems offer dynamic, context-aware suggestions.
-(2) LLM-Powered Explainable POI Recommendation.
+(2) LLMs-Powered Explainable POI Recommendation.
 Creating an explainable POI recommendation system via large language models is a promising avenue for leveraging advanced language processing capabilities to provide transparent and understandable POI recommendations.
@@
 This transparency enhances user trust and engagement by shedding light on the decision-making process behind each recommendation, ultimately improving the overall user experience of POI recommendation.
-Future POI recommendation architecture is to ensure that decentralized systems can deliver the same level of personalization and accuracy as their centralized counterparts, all while maintaining the privacy benefits that decentralized architectures provide.
-Small Language Model-Powered Decentralized POI Recommendation.
-While the mainstream focus is on developing large language models, small language models are also gaining attention for their efficiency and ability to run on-device, offering privacy-preserving and real-time capabilities.
-The newly released Llama 3.2’s lightweight, language models (1B and 3B)  [^1] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ can revolutionize edge AI and significantly improve the accuracy of decentralized POI recommendation systems by processing real-time, contextual data directly on the device.
-These models can learn from user interactions, preferences, and location data locally by using the long contexts (up to 128K tokens), allowing for highly personalized recommendations without needing centralized data aggregation.
-By running on-device, the models can continually adapt to user behavior, leading to more precise recommendations based on immediate context, while preserving privacy.
-The emergence of new models and architectures has brought up new opportunities, consequently expanding the landscape of potential threats.
-Future POI recommendation systems need to handle new security issues.
-(1) Addressing Vulnerabilities in Emerging POI Recommender Systems.
-As novel techniques such as multi-modal data infusion, self-supervised learning, and LLM-based recommendation systems gain traction, they introduce new layers of complexity and new vulnerabilities.
-The security landscape of POI recommendation
-systems will need to focus on understanding these emerging attack surfaces, as they differ significantly from traditional techniques.
-Future research should prioritize identifying and mitigating vulnerabilities unique to these advanced techniques, such as adversarial manipulation in multi-modal inputs or exploiting context-driven weaknesses in LLM-based systems.
-By developing robust security frameworks tailored to these innovative approaches, future POI recommendation systems can safeguard against both known and unforeseen threats, ensuring trust and reliability as these technologies evolve.
-(2) Privacy-Enhancing POI Recommendation based on Cryptography Techniques.
-Developing privacy-enhancing techniques for POI recommendation systems is crucial for protecting user data while maintaining recommendation quality.
-This involves implementing methods such as secure multiparty computation and zero-knowledge proof, ensuring user interactions and preferences remain confidential.
-By safeguarding sensitive information, these techniques can enhance user trust and encourage broader adoption of recommendation systems, thereby balancing privacy protection with system performance and ensuring that privacy-enhanced models do not compromise the accuracy and relevance of recommendations.
-By embracing these diverse research directions and leveraging cutting-edge technologies and methodologies, the field of POI recommendation is poised to revolutionize the way users discover and engage with points of interest in the digital landscape.
+(3) Semantic Alignment Between POI Models and Real Applications.
+Bridging semantic gaps between spatial data and real-world contexts enables POI models to understand POI spatial information.
+Integrating geographic hierarchies, proximity patterns, and spatial correlations among POIs (e.g., analyzing a park’s relation to transit hubs and neighborhoods) will improve POI recommendation contextual relevance, thereby improving POI recommendation accuracy.
+VII-B Architectures
+Future POI recommendation architectures will primarily be distributed, incorporating technologies such as decentralized and federated recommendation.
+In actual deployment, these approaches face efficiency bottlenecks, making it difficult to effectively meet the demands of large-scale systems.
+The following outlines several future research directions focusing on scalability, latency, and communication efficiency.
+(1) Scalability Enhancement in Distributed POI Recommendation.
+Improving the scalability of POI recommendation architectures is crucial for their effective deployment in large-scale systems.
+Current distributed systems struggle to accommodate an increasing number of computational nodes due to the overhead of large-scale data processing and inter-node coordination.
+The primary challenge lies in the need for efficient model training and result aggregation while maintaining computational efficiency.
+Future research could explore highly collaborative recommendation frameworks that dynamically balance computational workloads, optimize resource allocation, and simplify coordination among nodes to improve scalability and ensure efficient system expansion.
+(2) Latency Optimization for Real-Time POI Recommendation.
+Reducing latency is a fundamental requirement for achieving real-time POI recommendations, particularly in decentralized collaborative systems.
+These architectures rely on multiple distributed nodes to perform recommendation tasks, but frequent data synchronization, inter-node communication, and model parameter exchanges introduce significant delays, impacting responsiveness.
+To mitigate these issues, future research could focus on developing asynchronous communication protocols, parallel computation strategies, and lightweight update mechanisms to enhance real-time performance in decentralized POI recommendation systems.
+(3) Communication-Efficient Federated POI Recommendation.
+Federated recommendation models face significant challenges in handling the transmission of large-scale model parameters across distributed nodes [129, 131].
+As model sizes increase, resource-constrained nodes may struggle to efficiently participate in global model training, limiting the scalability of federated learning.
+To address this challenge, future research could focus on designing communication-efficient aggregation algorithms, adaptive parameter compression techniques, and incentive-driven task distribution mechanisms to optimize the efficiency of federated POI recommendation and enable seamless collaboration across diverse computational environments.
+VII-C Security
+Future research needs to address the challenge of balancing security and efficiency in existing privacy-preserving POI recommendation systems, as traditional techniques such as federated learning, differential privacy, and cryptographic methods often encounter efficiency issues due to their high computational or communication costs.
+The following research directions are worth exploring.
+(1) Application-Oriented Privacy-Preserving POI Recommendation.
+Optimizing privacy-preserving POI recommendation systems for specific application scenarios is crucial for improving efficiency.
+Different privacy protection methods exhibit varying trade-offs between security and computational cost.
+Federated learning excels in distributed training environments by safeguarding user privacy while minimizing data transmission overhead.
+Cryptographic techniques, such as homomorphic encryption and secure multi-party computation, are better suited for small-scale inference tasks that require strong privacy guarantees but lower computational complexity.
+Additionally, emerging machine unlearning techniques [173] offer efficient solutions for large-scale training and inference tasks, making them promising candidates for privacy-preserving POI recommendation.
+Future research should focus on designing adaptive frameworks that dynamically select the most suitable privacy-preserving strategy based on task characteristics such as training mode, inference complexity, and data scale.
+(2) TEE-Based Privacy-Preserving POI Recommendation.
+Leveraging Trusted Execution Environments (TEEs) presents a promising direction for enhancing security while maintaining computational efficiency in POI recommendation.
+TEEs, such as Intel SGX [174] and Arm TrustZone [175], provide secure enclaves for executing privacy-sensitive computations with minimal performance overhead.
+However, they remain susceptible to side-channel attacks and hardware vulnerabilities.
+Future research should explore techniques to enhance TEE security, such as mitigating known attack vectors and integrating lightweight cryptographic enhancements.
+By improving the robustness of TEEs, privacy-preserving POI recommendation systems can achieve secure and efficient computation without significant trade-offs in performance.
+Additionally, emerging cross-domain technologies, such as data watermarking, could be integrated to further enhance both security and efficiency in privacy-preserving POI recommendation systems.
+By combining technological customization with hardware optimization, future systems can improve privacy protection effectiveness while maintaining efficiency.
 VIII Conclusion